### v2

# Robust Online Learning with Private Information

Abstract
This paper investigates the robustness of online learning algorithms when learners possess private information. No-external-regret algorithms, prevalent in machine learning, are vulnerable to strategic manipulation, allowing an adaptive opponent to extract full surplus. Even standard no-weak-external-regret algorithms, designed for optimal learning in stationary environments, exhibit similar vulnerabilities. This raises a fundamental question: can a learner simultaneously prevent full surplus extraction by adaptive opponents while maintaining optimal performance in well-behaved environments? To address this, we model the problem as a two-player repeated game, where the learner with private information plays against the environment, facing ambiguity about the environment’s types: stationary or adaptive. We introduce partial safety as a key design criterion for online learning algorithms to prevent full surplus extraction. We then propose the Explore-Exploit-Punish (EEP) algorithm and prove that it satisfies partial safety while achieving optimal learning in stationary environments, and has a variant that delivers improved welfare performance. Our findings highlight the risks of applying standard online learning algorithms in strategic settings with adverse selection. We advocate for a shift toward online learning algorithms that explicitly incorporate safeguards against strategic manipulation while ensuring strong learning performance.
1 Introduction
The increasing reliance on online platforms and algorithmic decision-making has underscored the importance of agents’ ability to learn and adapt in complex, dynamic environments. In settings such as online advertising auctions and personalized pricing, economic agents engage in repeated interactions where learning to optimize strategies is crucial. A dominant approach in these contexts is online learning, where algorithms iteratively refine their decisions based on observed outcomes. However, many commonly studied design objectives assume a stationary or non-strategic environment. In economically relevant settings—particularly those involving strategic interactions and private information—this assumption often fails, exposing learning agents to significant vulnerabilities. This raises a fundamental question: what constitutes a well-designed online learning algorithm in strategic environments?
The central contribution of this paper is to identify the vulnerability of standard online learning algorithms in strategic environments, introduce a new design objective, and propose an algorithm that satisfies it. We first show that many standard online learning algorithms are unsafe when facing an adaptive and strategic opponent, meaning that such an opponent can exploit the learning dynamics of algorithms to extract nearly the entire surplus from a privately informed learner. This finding raises serious concerns about the direct applicability of off-the-shelf online learning algorithms in economic settings with adverse selection. To design online learning algorithms that learners with private information can safely use in strategic settings, we introduce partial safety as a key desideratum for online learning algorithms. Partial safety is a conservative notion of robustness aimed at preventing full surplus extraction by adaptive opponents. Since standard online learning algorithms are not partially safe, it remains unclear ex-ante whether an algorithm can be designed to ensure both partial safety and effective learning in certain environments. We construct an example of an online learning algorithm that is partially safe and can achieve optimal learning in stationary environments.
The model we analyze is a two-player repeated game played by a learner and an environment. In each period, the learner selects an action, while the environment chooses a mechanism—an allocation and payment rule. The learner has a private type that is unknown to the environment before the interaction begins. The learner’s payoff depends on an allocation, payment, and their private type; the environment’s payoff only depends on an allocation and payment. There are two possible types of environments: a stationary environment, which follows a fixed mixed strategy in each period, and an adaptive environment, which adaptively responds to the learner’s online learning algorithm to maximize its payoff. Before play begins, the learner selects an online learning algorithm, which determines a strategy given the learner’s private type, while facing ambiguity about the environment’s type.
Given this ambiguity in the game structure, the literature lacks a standard behavioral assumption that defines rationality on the learner’s side. This raises an important question: What constitutes a “good” learning algorithm in strategic settings? We propose that a desirable design objective is to balance two competing considerations: (i) achieving optimal performance in favorable settings, such as stationary environments, while (ii) protecting against strategic manipulation in adversarial settings, such as adaptive environments. In other words, an effective learning algorithm should hope for the best—exploiting opportunities for efficient learning when conditions allow—while preparing for the worst—safeguarding against worst-case exploitation. Our analysis investigates whether and how online learning algorithms can achieve both objectives simultaneously.
Our analysis yields several key results. First, we show that no-external-regret (no-ER) algorithms, while effective in stationary settings, are inherently unsafe when facing an adaptive environment. The no-ER condition requires that if the environment repeatedly selects the same action over an extended period, the learning algorithm must eventually reveal a best response to that action. An adaptive environment can exploit this property by strategically probing the learner’s responses in early rounds to infer their private type, enabling it to extract the learner’s surplus in later stages of the interaction.
Second, we show that several standard no-weak-external-regret (no-WER) algorithms—designed to learn optimal actions in stationary environments—are also unsafe. In particular, we demonstrate that Uniform Exploration (UE), Successive Elimination (SE), and UCB are unsafe. UE and SE are unsafe because they eliminate all but one action during the exploration phase, allowing an adaptive environment to simply wait until exploration concludes before extracting surplus. For UCB, we construct an adaptive environment’s strategy that exploits the algorithm’s adaptive behavior to infer the learner’s type. These findings cast doubt on the applicability of off-the-shelf online learning algorithms in strategic settings where the learner possesses private information.
Third, we introduce an online learning algorithm, Explore-Exploit-Punish (EEP), that satisfies no-weak-external-regret (no-WER) while ensuring partial safety. EEP builds upon Uniform Exploration but incorporates a mechanism to guard against strategic manipulation when the environment is adaptive. The algorithm operates in three distinct phases. In the exploration phase, EEP selects each action in a round-robin manner for a pre-specified number of periods, constructing confidence intervals for each action’s expected allocation and payment. These confidence intervals are designed to contain the true expected values if the environment is stationary. In the exploitation phase, the algorithm selects the action that demonstrated the highest performance during exploration and forms a new set of confidence intervals for this action. If these new confidence intervals are inconsistent with those established in the exploration phase, the algorithm transitions to the punishment phase, in which it ceases participation to deter further exploitation. This strategic use of confidence intervals makes it difficult for an adaptive environment to manipulate the learner while ensuring that punishment is unlikely to be triggered in a stationary environment. As a result, EEP maintains strong learning performance under no-WER while offering protection against adaptive opponents. Furthermore, we show that a variant of EEP can achieve welfare efficiency while maintaining no-WER and partial safety.
These findings call for a re-evaluation of the design principles for online learning algorithms in strategic economic settings. We advocate for a shift toward algorithms that explicitly prioritize robustness against strategic manipulation while ensuring strong learning performance in well-specified environments, rather than optimizing a generic regret measure without specifying the relevant class of environments.
The paper proceeds as follows. Section 2 introduces the model. Section 3 presents our main results, including the vulnerabilities of existing algorithms and the properties of EEP and ESEP. Section 5 concludes with a discussion of future research directions.
1.1 Related work
This paper builds upon and contributes to several strands of literature.
Exploiting no-regret-learning agents
A growing body of work examines how a principal (mechanism designer) can exploit agents using no-regret learning algorithms in repeated games with adverse selection. Braverman et al. (2017) initiated this line of research, showing that an auction designer can extract the full surplus from an agent if the agent’s learning algorithm is mean-based, a subclass of no-external-regret (no-ER) algorithms. Deng et al. (2019a) extended this result to prior-free settings, while Kumar et al. (2024) analyzed online gradient descent (OGD) algorithms, demonstrating their strategic robustness as they prevent the seller from earning more than the Myerson-optimal revenue. Guruganesh et al. (2024) explored a related setting involving moral hazard.
This paper also studies online learning algorithms with adverse selection but differs in several key respects. First, unlike the previous studies, which allow the agent’s type to evolve as long as it is independently drawn from a fixed distribution and observed at the beginning of each period, we consider a setting where the agent’s private type remains fixed. This enables us to establish a general impossibility result for no-ER algorithms without relying on the mean-based assumption (Theorem 1.) Furthermore, rather than focusing solely on impossibility results concerning safety, we aim to identify desirable design objectives from the learner’s perspective in light of these limitations.
Learning algorithms in strategic environment
Game theorists have long studied no-regret algorithms for their role in providing microfoundations for equilibrium concepts in static games (Hart and Mas-Colell, 2000, 2001). More recent work examines the payoff guarantees of these algorithms and the incentives for agents to adopt them. Deng et al. (2019b) and Mansour et al. (2022) show that when a principal interacts with a learning agent using a no-swap-regret algorithm, the principal’s average payoff cannot exceed the Stackelberg value. Meanwhile, Arunachaleswaran et al. (2024) analyze settings where a learner strategically selects an online learning algorithm against an optimizer who maximizes her own payoff given the learner’s choice. They show that no-swap-regret algorithms are Pareto-optimal, meaning no alternative algorithm achieves strictly better average payoffs for every possible optimizer’s payoff function.
These studies do not explicitly account for the role of private information in learning. While no-swap-regret algorithms possess certain desirable incentive properties, they remain a subclass of no-ER algorithms. Consequently, our impossibility result implies that even no-swap-regret algorithms are susceptible to full surplus extraction by a strategic opponent. Unlike Arunachaleswaran et al. (2024), which considers all possible payoff functions of the optimizer, we focus on a specific class of principal payoff functions that naturally arise in economic settings.
Repeated games
This paper also relates to the literature on repeated games, particularly those with incomplete information (see Renault (2020) for a survey). Hart (1985) examines a general class of repeated games with incomplete information, while Shalev (1994) focuses on a more tractable yet still rich subclass known as repeated games with known payoffs, which closely resembles our setting. These studies assume that players are strategically sophisticated, possess knowledge of the opponent’s strategy, and best respond accordingly. In contrast, this paper—along with other work on learning in games—focuses on agents who lack strategic sophistication and instead rely on learning algorithms to perform well.
2 Model
2.1 Setup
A decision-maker called a learner (agent, he) plays a two-player $T$-period repeated game against an environment (principal, she). The players commonly know the finite horizon $T\in\mathbb{Z}_{>0}$. Let $\mathcal{A}$ denote a finite stage-game action set (e.g., the set of bids) of the learner with $|\mathcal{A}|\geq 2$. Let $\mathcal{M}\coloneqq\{0,1\}^{\mathcal{A}}\times[0,1/2]^{\mathcal{A}}$ denote the set of the environment’s stage-game actions. A typical element $(x,p)\in\mathcal{M}$ is a mechanism, or a pair of an allocation rule $x\colon\mathcal{A}\to\{0,1\}$ and a payment rule $p\colon\mathcal{A}\to[0,1/2]$. [^2] The upper bound $1/2$ is just chosen for normalization. All the argument goes through as long as the range is bounded.
Before the start of play, the learner’s type $\theta\in\Theta$ is chosen once and for all, where $\Theta\subseteq[0,1/2)$ is a commonly known finite set. In each period, the payoffs are determined as follows: the learner and the environment simultaneously choose their actions $a\in\mathcal{A}$ and $(x,p)\in\mathcal{M}$. The learner’s stage-game payoff is $u(a,(x,p),\theta)\coloneqq\theta x(a)-p(a)$. The learner can always opt-out if he wants: there exists a special action $a_{0}\in\mathcal{A}$ such that $x(a_{0})=0$ and $p(a_{0})=0$ for any $(x,p)\in\mathcal{M}$. The players care about undiscounted time-averaged payoffs. [^3] The environment’s stage-game payoff does not depend on the agent’s type. This setup is called learning with known own payoffs (cf. Shalev (1994).)
A (behavioral) strategy of a player is a mapping from each player’s private history up to the previous period to a distribution over actions for the current period. A player’s private history is composed of their past observations, though the information each player receives differs based on the components of the game they observe. For the learner, at the end of each period $t$, he observes the allocation $x_{t}(a_{t})$ and the payment $p_{t}(a_{t})$ given the stage game action profile $(a_{t},(x_{t},p_{t}))$. Formally, a private history of the learner up to time $t$ is defined as $h^{t}_{A}\coloneqq\left(\theta,\left(a_{s},x_{s}(a_{s}),p_{s}(a_{s})\right)_{s=1}^{t}\right)$. Notably, the learner does not have full information regarding the environment’s action $(x_{t},p_{t})$ in the sense that he does not observe the value of $(x_{t}(a),p_{t}(a))$ for $a\neq a_{t}$.  [^4] The learner’s problem is a specific instance of the class of problems known as learning with partial feedback, or bandit feedback, in the statistical learning literature. This setting is closely related to repeated games with imperfect monitoring (Lehrer and Solan (2016)) and repeated games with incomplete information and private learning (Wiseman (2012)) in game theory. In contrast, the environment observes both the learner’s actions and its own mechanism but does not observe the agent’s type: a private history of the environment up to time $t$ is $h^{t}_{P}\coloneqq(a_{s},x_{s},p_{s})_{s=1}^{t}$.
An online learning algorithm outputs a behavioral strategy of the learner for any given agent’s type $\theta$ and time horizon $T$. Formally, online learning algorithms are defined as follows:
Definition 1 (Online learning algorithms).
Let $\mathcal{H}_{A}^{T}$ denote the set of learner’s private histories with length less than $T$ and denote the set of behavioral strategies for a $T$-period repeated game by $\mathcal{S}_{A}^{T}\coloneqq(\Delta(\mathcal{A}))^{\mathcal{H}_{A}^{T}}$. [^5] For a topological space $A$, the set of all Borel probability measures on $A$ is denoted by $\Delta(A)$. An online learning algorithm $\mathcal{L}$ is a mapping from $(\theta,T)\in\Theta\times\mathbb{Z}_{>0}$ to $\sigma_{A}^{T}\in\mathcal{S}_{\mathcal{A}}^{T}$.
The learner interacts with a potentially adaptive environment, which may belong to one of two possible types. An adaptive-type environment observes the learner’s chosen online learning algorithm (without knowing the agent’s private type $\theta$) and subsequently selects a behavioral strategy $\sigma_{P}^{T}\in\mathcal{S}_{P}^{T}$, where $\mathcal{S}_{P}^{T}$ denotes the set of behavioral strategies available to the environment. [^6] Even if the environment is aware of the agent’s online learning algorithm, it may still have uncertainty about the agent’s type. The adaptive-type environment’s stage-game payoff is given by $v(a,(x,p))\coloneqq p(a)$, and it aims to maximize its undiscounted time-averaged payoff. We assume that an adaptive-type environment is Bayesian, holding a belief with full support $\pi\in\Delta(\Theta)$ over the type space $\Theta$. We denote such an environment as $A(\pi)$. The belief $\pi$ is unknown to the learner.
In contrast, a stationary-type environment commits to a fixed mixed strategy in each stage game. A mixed strategy specifies a probability distribution over allocations and payments for each stage-game action of the learner. Formally, the set of mixed strategies is given by $\overline{\mathcal{S}}_{P}\coloneqq\Delta\left(\left(\{0,1\}\times[0,1/2]\right)^{\mathcal{A}}\right).$
The learner, facing ambiguity about the environment’s type, employs an online learning algorithm to select strategies that perform “well” across both types of environments. [^7] Rigorously speaking, there are two possible classes of types. The class of adaptive types is parametrized by belief $\pi$; the class of stationary types is parametrized by mixed strategies. We formalize this notion in subsequent sections.
Example 1 (Online advertisement auction).
Advertisers in online advertising auctions use bidding algorithms to optimize decisions over time while facing ambiguity about the auction environment (e.g., the number of competing bidders or the details of the mechanism). [^8] For institutional details, see Choi et al. (2020); Despotakis et al. (2021); Tunuguntla and Hoban (2021); Zeithammer (2019), among others. The action set $\mathcal{A}$ represents possible bids. For each customer segment (e.g., users with a specific query), an advertiser selects a bid $a_{t}$ and observes the outcome $(x_{t}(a_{t}),p_{t}(a_{t}))$, where $x_{t}(a_{t})$ indicates whether the bid wins and $p_{t}(a_{t})$ is the corresponding payment. The advertiser’s type $\theta$ reflects their private valuation for securing an ad slot.
The auction environment is stationary if competing bidders’ bids follow a fixed distribution and the platform maintains stable reserve prices. Even if the platform updates its system periodically, the environment remains effectively stationary as long as these updates occur infrequently and are not heavily dependent on any single advertiser’s learning process. Conversely, the environment is adaptive and strategic if the auctioneer (e.g., Google or Amazon) frequently adjusts reserve prices to maximize revenue (Zeithammer, 2019), particularly when an advertiser effectively becomes the sole bidder for certain customer segments due to sophisticated targeting mechanisms. Such strategic behavior is especially relevant when auctioneers leverage detailed user-level data for price discrimination. [^9] For instance, a retargeting advertiser—whose website the customer recently visited before entering the auction—likely values the impression far more than advertisers bidding only on demographics.
A potentially adaptive environment arises when the advertiser is uncertain about the type of environment they face in an auction. This uncertainty may stem from a lack of transparency in the auction rules or from limited information about how the platform adapts its pricing strategies over time.
2.2 Popular design goals in machine learning literature
In the current setup, the learner does not exactly know about the environment’s type and strategy. We also assume that he has no prior over these objects, possibly due to a lack of past data and/or experience. Given such ambiguity in the game structure, the literature does not seem to have a standard behavioral assumption that captures the rationality on the learner’s side. [^10] Camara et al. (2020) and Collina et al. (2023) introduce a regret notion to capture learners’ rationality in prior-free settings, where both a principal and an agent simultaneously learn the state distribution. Unlike our model, the agent in their framework does not have private types.
Satisfying a condition called no external regret (no-ER) is a popular design goal of online learning algorithms in the literature of machine learning. The external regret of a strategy against sequences of types and environment’s actions is defined as the difference between the payoff of the best-fixed action in hindsight and the expected payoff of the actions chosen by the strategy, fixing the environment’s actions. The no-ER condition requires the algorithm to achieve, in expectation, an average payoff at least as good as the one achieved by the best-fixed action in hindsight for any realized action and type sequences.
Definition 2 (No external regret (No-ER)).
[…]
$\mathrm{ER}\left(\sigma_{A}^{T};T,\theta,(x_{1:T},p_{1:T})\right)\coloneqq\max_{a\in\mathcal{A}}\sum_{t=1}^{T}u(a,(x_{t},p_{t}),\theta)-\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T}u(a_{t},(x_{t},p_{t}),\theta)\right],$ (1)
where the expectation is taken over the distribution of learner’s action sequences $a_{1:T}$ induced by $\sigma_{A}^{T}$ and $(x_{1:T},p_{1:T})$. We say an online learning algorithm $\mathcal{L}$ has no external regret (no-ER) if there exists $R\colon\mathbb{Z}_{>0}\to\mathbb{R}_{\geq 0}$ such that for any $T$, $\theta$, and $(x_{1:T},p_{1:T})$, we have
$\mathrm{ER}\left(\mathcal{L}(\theta,T);T,\theta,(x_{1:T},p_{1:T})\right)\leq R(T),$ (2)
with $R(T)=o(T)$.121212 For function $R\colon\mathbb{Z}_{>0}\to\mathbb{R}_{\geq 0}$, $R(T)=o(T)$ means $\limsup_{T\to\infty}\frac{R(T)}{T}=0$.
We often show the regret upper bound in the form $R(T)=C\left(\log T\right)^{\delta}T^{\gamma}$ for some constants $C>0$, $\delta\in[0,1)$ and $\gamma\in(0,1)$ that are independent of $T$. Note that if $\gamma<1$, regardless of $C$ and $\gamma$, we have $R(T)=o(T)$.
Definition 3 ($(C,\delta,\gamma)$-no-ER).
Let $C>0$, $\delta\in[0,1)$ and $\gamma\in(0,1)$. We say a no-ER algorithm $\mathcal{L}$ has $(C,\delta,\gamma)$-no-ER if $\mathrm{ER}\left(\mathcal{L}(\theta,T);T,\theta,(x_{1:T},p_{1:T})\right)\leq C\left(\log T\right)^{\delta}T^{\gamma}$ for any $T$, $\theta$ and $(x_{1:T},p_{1:T})$. We say an algorithm has $\gamma$-no-ER if it has $(C,\delta,\gamma)$-no-ER for some $C>0$ and $\delta\in[0,1)$.
Below is a well-known example of a no-ER algorithm.
Example 2 (EXP3).
EXP3 (Algorithm 2 in Appendix A.4) has $\left(\sqrt{2|\mathcal{A}|\log|\mathcal{A}|},0,\frac{1}{2}\right)$-no-ER. [^13] See Theorem 11.2 of Lattimore and Szepesvári (2020) for a textbook reference. The algorithm puts more weight on the actions that historically perform well. When updating the weights, instead of using the actual observed reward, it uses the unbiased estimator of the reward.
Another possible design goal of online learning algorithms is to achieve a vanishing external regret only for stationary distributions.
Definition 4 (No weak external regret (No-WER)).
Given horizon $T$, agent’s type $\theta$, and environment’s mixed strategy $\sigma_{P}\in\overline{\mathcal{S}}_{P}$, the weak external regret of learner’s strategy $\sigma_{A}^{T}$ is defined as
$\mathrm{WER}\left(\sigma_{A}^{T};T,\theta,\sigma_{P}\right)\coloneqq\sum_{t=1}^{T}\max_{a\in\mathcal{A}}\mathop{\mathbbm{E}}_{\sigma_{P}}\left[u(a,(x_{t},p_{t}),\theta)\right]-\mathop{\mathbbm{E}}_{\sigma_{A}^{T},\sigma_{P}}\left[\sum_{t=1}^{T}u(a_{t},(x_{t},p_{t}),\theta)\right],$ (3)
where the expectation is taken over the distribution of a tuple of actions $(a_{1:T},(x_{1:T},p_{1:T}))$ induced by $\sigma_{A}^{T}$ and $\sigma_{P}$. We say an online learning algorithm $\mathcal{L}$ has no weak external regret (no-WER) if there exists $R\colon\mathbb{Z}_{>0}\to\mathbb{R}_{\geq 0}$ such that for any $T$, $\theta$, and $\sigma_{P}\in\overline{\mathcal{S}}_{P}$,
$\mathrm{WER}\left(\mathcal{L}(\theta,T);T,\theta,\sigma_{P}\right)\leq R(T),$ (4)
with $R(T)=o(T)$. The concepts of $(C,\delta,\gamma)$-no-WER and $\gamma$-no-WER are defined similarly to no-ER.
Remark 1.
Since $\sigma_{P}$ is a fixed mixed strategy, there is one action that is optimal against $\sigma_{P}$ throughout all periods. Thus, the first term of the LHS of (3) is equal to $\max_{a\in\mathcal{A}}\sum_{t=1}^{T}\mathop{\mathbbm{E}}_{\sigma_{P}}\left[u(a,(x_{t},p_{t}),\theta)\right]$.
Although many studies in the literature on learning in games analyze the dynamics of games assuming that agents are using no-ER algorithms, no-WER is also used as a design goal in adjacent fields such as stochastic multi-armed bandit. Below are examples of standard no-WER algorithms.
Example 3 (Uniform Exploration).
Uniform Exploration is the one of the simplest no-WER algorithms. It tries each arm $T_{1}$ times, and then chooses the arm with the highest average reward in all remaining rounds. It is well-known that if we choose $T_{1}\coloneqq T^{2/3}(\log T)^{1/3}$, suitably to balance the exploration-exploitation trade-off, then Uniform Exploration has $\frac{2}{3}$-no-WER. [^14] See Chapter 1 of Slivkins (2019) for a textbook reference to well-known no-WER algorithms.
Example 4 (Successive Elimination).
Successive Elimination is a $\frac{1}{2}$-no-WER algorithm. It eliminates actions once they are estimated to be suboptimal (see Algorithm 3 in Appendix A.4.) To estimate the performance of each action, the algorithm constructs a confidence interval for the expected payoff of each action. It eliminates actions whose upper confidence bound (UCB) is below the lower confidence bound (LCB) of any other action. If the environment is stationary, the confidence interval contains the true mean $\mathbbm{E}_{\sigma_{P}}[u(a,(x,p),\theta)]$ for all $t$ and $a$ with high probability (see Appendix A.2.)
Example 5 (UCB).
UCB, which stands for upper confidence bound, is another $\frac{1}{2}$-no-WER algorithm. It balances exploration and exploitation by choosing the action with the highest upper confidence bound in each period (see Algorithm 4 in Appendix A.4.)
2.3 Other possible design goals in potentially adaptive environments
Due to its applicability, more and more people use off-the-shelf online learning algorithms in many different domains, including some potentially adaptive environments like online ad auctions. Indeed, there is empirical evidence suggesting firms use no external regret algorithms in online ad auctions (Nekipelov et al., 2015). Given this trend, one might seek online learning algorithms that are “robust” to use in such environments. Below, we introduce a conservative notion of robustness: an algorithm is partially safe if no environment’s strategy can consistently extract almost all of the surplus in the long run.
Definition 5 (Unsafe/partially safe).
We say an online learning algorithm $\mathcal{L}$ is unsafe if,
$\forall\varepsilon>0\ \exists\overline{T}\ \forall T\geq\overline{T},\ \exists\sigma_{P}^{T}\ \forall\theta\in\Theta,\quad\mathop{\mathbbm{E}}_{\sigma_{A}^{T},\sigma_{P}^{T}}\left[\frac{1}{T}\sum_{t=1}^{T}v\bigl{(}a_{t},(x_{t},p_{t})\bigr{)}\right]\geq\theta-\varepsilon,$ (5)
where the expectation is taken over $(\mathcal{A}\times\mathcal{M})^{T}$ with respect to the distribution induced by $\sigma_{A}^{T}\coloneqq\mathcal{L}(\theta,T)$ and $\sigma_{P}^{T}$. We say an online algorithm $\mathcal{L}$ is partially safe if it is not unsafe.
If the learner employs an unsafe algorithm, then for sufficiently large $T$, the learner’s payoff will always be approximately zero when facing an adaptive-type environment. As shown in Section 3, many existing online learning algorithms fail to satisfy this conservative robustness criterion.
While ensuring partial safety is preferable to having no safeguard at all, it may still be insufficient. For instance, an agent that always selects $a_{0}$ is partially safe since this guarantees that the environment’s payoff is zero. However, this is clearly inefficient in terms of welfare. This motivates another desirable property for online learning algorithms:
Definition 6 (Welfare efficient).
Suppose that the environment is adaptive. We say an online learning algorithm $\mathcal{L}$ is welfare efficient if, for any $\varepsilon>0$, there exists $\overline{T}$ such that for any $T\geq\overline{T}$, $\theta\in\Theta$, and $\pi\in\Delta(\Theta)$, the sum of ex-ante expected payoffs of the learner and the environment is $\theta-\varepsilon$ when agent’s type is $\theta$ and $A(\pi)$ best-responds. [^15] The expectation is taken over action sequences induced by the strategies of the agent and the adaptive environment. Note that $\theta$ and $\pi$ are fixed.
Whether an online learning algorithm can simultaneously satisfy no-WER, partial safety, and welfare efficiency is not evident ex-ante. In Section 3, we provide an affirmative answer to this question.
3 Results
3.1 Preliminary results
As its name suggests, no-WER is a weaker requirement than no-ER.
Lemma 1.
Any no-ER algorithm has no-WER.
Proof.
See Appendix B.3
3.2 Main results
Figure 1 provides an overview of the main results presented in this section. Section 3.2.1 establishes that many existing online learning algorithms are unsafe. Notably, any unsafe algorithm must be welfare-efficient, as the adaptive environment can extract the full surplus. This explains why there is no algorithm that is both unsafe and welfare-inefficient while also being no-WER. Section 3.2.2 introduces EEP, an algorithm that has no-WER and partially safe. Section 3.2.3 further examines its variant, ESEP, which has no-WER, partially safe, and welfare-efficient. Importantly, even highly unsophisticated algorithms can satisfy some of these design criteria. For instance, an algorithm that always selects $a_{0}$ is partially safe, whereas one that consistently chooses $\overline{a}\neq a_{0}$ allows the adaptive environment to extract full surplus, making it welfare-efficient. Lastly, we can show that the equilibrium strategies proposed in repeated game (RG) literature are partially safe and welfare efficient, but they are not no-WER since such strategies assume that both players are strategically sophisticated and best respond to each other.
Figure 1: Summary of Results
3.2.1 Unsafeness of existing algorithms
The No-ER condition (Definition 2) mandates that the agent reveals their optimal response to any fixed action $(x,p)\in\mathcal{M}$ taken by the principal if the principal consistently selects $(x,p)$ for a sufficiently long period. Leveraging this characteristic, the principal can swiftly ascertain the agent’s type in the initial phase of the game and subsequently exploit this information for the remainder of the interaction.
Theorem 1.
Any no-ER algorithm is unsafe.
Proof.
See Appendix B.4. ∎
Remark 2.
The proof of Theorem 1 shows that the environment can extract full surplus even when it is restricted to a standard auction in which $p_{t}(a)>0$ only when $x_{t}(a)=1$.
Remark 3.
Theorem 6 of Brown et al. (2023) implies the same conclusion as Theorem 1, with an additional assumption that no-ER condition is satisfied anytime unlike Theorem 1 (see Brown et al. (2023) for the definition of anytime no-ER algorithms.) The principal-agent setup satisfies their Assumption 2, so by Theorem 6, for any $\varepsilon>0$, there exists $T_{0}$ such that the strategic environment finds an $\varepsilon$-approximate Stackelberg strategy $(x_{0},p_{0})\in\mathcal{M}$ in period $T_{0}$. If $T$ is sufficiently long, then if $(x_{0},p_{0})$ is chosen for all periods after $T_{0}$, the average payoff of the principal would be $V(\theta)-\varepsilon+o(1)$.
Theorem 1 does not exclude the possibility that there exists some no-WER algorithm that is partially safe. However, standard no-WER algorithms in the literature are shown to be unsafe as well.
Proposition 1.
  1. 1.
    Uniform Exploration and Successive Elimination are no-WER and unsafe.
  2. 2.
    UCB has no-WER. Moreover, if $|\mathcal{A}|\geq|\Theta|+1$, then UCB is unsafe.
Proof.
See Appendix B.1. ∎
Remark 4.
The condition $|\mathcal{A}|\geq|\Theta|+1$ is relevant when the bidding space is sufficiently large, which is often the case in real-world online advertisement auctions. Note that we did not show that this condition is necessary to establish that UCB is unsafe.
3.2.2 Existence of algorithms that are partially safe and of no-WER
We show the existence of a no-WER algorithm that ensures partial safety, referred to as the Explore-Exploit-Punish (EEP) (Algorithm 6). The EEP operates in three distinct phases. During Phase 1, the algorithm selects actions round-robin for pre-fixed periods $|\mathcal{A}|T_{1}$, establishing confidence intervals for each action. At the end of Phase 1, the algorithm chooses the action with the highest empirical reward, recording the confidence bounds of the allocation and payment for the chosen action, and transitions to the second phase. In Phase 2, the algorithm forms new confidence intervals for the allocation and payment. Transition to the third phase occurs when the new confidence bounds established in Phase 2 are found to be inconsistent with the confidence bounds established in Phase 1. In Phase 3, the algorithm consistently plays the default action $a_{0}$. Theorem 2 states that the EEP is both no-WER and partially safe, with a suitable choice of exploration length $T_{1}$ balancing the exploration-exploitation trade-off. Here, we provide a sketch of the proof.
For the no-WER property, when the environment is stationary, Phase 1 of the EEP algorithm resembles Uniform Exploration, identifying the optimal action with high probability. Once the best action is identified and the environment remains stationary, the algorithm is unlikely to trigger Phase 3, meaning its overall performance closely mirrors that of Uniform Exploration, which is known to possess the no-WER property. Therefore, EEP inherits the no-WER guarantee.
In terms of partial safety, consider that in order for the adaptive environment to achieve any time-average payoff $\theta$ for $\theta\in\Theta$, the durations of Phases 1 and 3—where the environment incurs strictly positive losses on average—must be of order $o(T)$, regardless of the learner’s type. To fully extract the surplus in Phase 2, the environment must ensure that the learner does not select action $a_{0}$, irrespective of his type $\theta$. This requires that, independent of $\theta$, the environment makes the learner choose an action $a\neq a_{0}$, providing a higher empirical mean payoff than 00. Roughly speaking, during Phase 1, the environment posts a price $p$, and the learner chooses $a\neq a_{0}$ if $\theta>p$, and thus we must have $p<\theta$ for all $\theta>0$. Since the true mean $\theta$ is higher than price $p$ posted in Phase 1, to achieve the average payoff $\theta$, the environment must increase the price more than $p$ during Phase 2. However, such an attempt by the environment would be quickly detected by the confidence intervals formed by the learning algorithm, thereby preventing prolonged exploitation. The following lemma supports such preventive use of confidence intervals in the EEP algorithm, claiming that the gain the environment can obtain via manipulation during Phase 2 is sublinear (see also its use in the proof of Theorem 2.)
Lemma 2 (Preventive use of confidence intervals).
For $s\in\mathbb{Z}_{>0}$, let $\overline{\rho}_{s}\coloneqq\sqrt{(2\log T)/s}$. Given a real sequence $(c_{t})_{t\geq 1}$, let $\overline{c}_{s}\coloneqq s^{-1}\sum_{t=1}^{s}c_{t}$.
Fix any $T_{0}\in\mathbb{Z}_{>0}$, $\Delta>0$, and $B\in\mathbb{R}$. Suppose that $\overline{c}_{s}+\overline{\rho}_{s}\geq B$ for all $s\leq T_{0}$ and $\overline{c}_{T_{0}}\leq B-\Delta$. Then, we have
$T_{0}\leq\frac{2\log T}{\Delta^{2}}.$ (6)
Similarly, if $\overline{c}_{s}-\overline{\rho}_{s}\leq B$ for all $s\leq T_{0}$ and $B+\Delta\leq\overline{c}_{T_{0}}$, then we have (6).
Proof.
See Appendix B.5. ∎
Theorem 2 (Properties of EEP).
EEP (Algorithm 6) with $T_{1}\coloneqq T^{2/3}\left(\log T\right)^{1/3}$ has no-WER and is partially safe.
Moreover, suppose that the environment type is $A(\pi)$. Define the monopoly price $p(\pi)$ by
$p(\pi)\coloneqq\sup\left\{\mathop{\rm arg~{}max}\limits_{p\in\Theta}\mathbbm{E}_{\pi}\left[p\mathbbm{1}\{\theta\geq p\}\right]\right\}.$
Then, for any sufficiently large $T$, conditional on $\theta$ being realized, the environment’s expected payoff is
$\mathbbm{1}\{\theta\geq p(\pi)\}p(\pi)+o(1);$
and the learner’s expected payoff is
$\mathbbm{1}\{\theta\geq p(\pi)\}\left(\theta-p(\pi)\right)+o(1).$
Proof.
See Appendix B.6. ∎
Remark 5.
The construction of EEP depends on the knowledge of $T$. There are common tricks to make algorithms anytime (i.e., knowledge of $T$ is not required to run the algorithms), and the doubling trick is one of them (see Appendix B.2.) Whether we can apply the doubling trick to make EEP anytime depends on the environment’s ability to track the identity of the learner. If the learner can restart his algorithm without revealing his identity to the environment, then we can apply the doubling trick to EEP and obtain an anytime algorithm that has no-WER and partially safe.
3.2.3 Welfare efficiency of EEP
EEP is not welfare efficient (Definition 6) because, intuitively, the adaptive environment can only post one price that is optimal under its prior. If the price posted during Phase 1 is not the right one and the learner chooses $a_{0}$ afterward, we experience welfare loss. Can we make an algorithm that satisfies all three desirable properties: partial safety, no-WER, and welfare efficiency?
To construct such an algorithm, we use one trick from the literature of repeated games. Since the type space $\Theta$ is finite, we can encode each $\theta\in\Theta$ as $|\mathcal{A}|$-ary number using actions as alphabets. As a result, the learner can spend $\lceil\log_{|\mathcal{A}|}(|\Theta|)\rceil$ periods to signal his type to the environment, and the effect of such a signaling phase on time-averaged payoff is negligible when $T$ is large.
Example 6.
Suppose that $\Theta\coloneqq\{\theta_{1},\theta_{2},\theta_{3}\}$ and $\mathcal{A}\coloneqq\{a_{0},a_{1}\}$. Then, the learner can signal any type by spending two periods with the following mapping:
$\begin{pmatrix}a_{0}a_{0}&a_{1}a_{0}&a_{1}a_{0}\\ \theta_{1}&\theta_{2}&\theta_{3}\end{pmatrix}.$
With this observation, we can consider the following variant of EEP:
  1. 1.
    Phase 1 (Exploration): this phase is the same as EEP. The estimated best action $a^{*}$ is chosen at the end of this phase, with the estimated payment: Let
    $\overline{p}\coloneqq\frac{1}{T_{1}}\sum_{t=1}^{|\mathcal{A}|T_{1}}\mathbbm{1}\{a_{t}=a^{*}\}p_{t}(a_{t}).$
  2. 2.
    Phase 2 (Signaling): Use $\lceil\log_{|\mathcal{A}|}(|\Theta|)\rceil$ periods to signal agent’s true type $\theta$.
  3. 3.
    Phase 3 (Exploitation with Protection):
    • if $a^{*}\neq a_{0}$, then this phase is the same as Phase 2 of EEP;
    • if $a^{*}=a_{0}$, then the algorithm chooses some pre-specified action $\overline{a}\neq a_{0}$ during this phase, and form the confidence interval in the same way as Phase 2 of EEP. The algorithm enters Phase 4 iff at least one of the following conditions is satisfied in some period $t$:
      $\mathrm{UCB}_{x}^{2}(t)<1,\quad\mathrm{LCB}_{p}^{2}(t)>\varepsilon_{p},$
      where $\varepsilon_{p}\in(0,\theta)$ is some pre-specified number.
  4. 4.
    Phase 4 (Punishment): same as Phase 3 of EEP.
We call this strategy ESEP (Explore-Signal-Exploit-Punish). Theorem 3 shows that ESEP has no-WER, is partially safe, and achieves welfare efficiency. Below is a concise sketch of the proof.
First, we can show that, by construction of ESEP, the analysis of the game essentially boils down to studying the following dynamic game, which captures the gameplay on the clean event (i.e., the event that happens with probability $1-o(1)$):
  1. 1.
    The agent privately observes his type $\theta$;
  2. 2.
    The environment offers a price $p$;
  3. 3.
    The agent signals his type $\widetilde{\theta}$;
  4. 4.
    The environment offers price $p^{\prime}$, which could be different from $p$;
  5. 5.
    The agent chooses action $a$. Payoffs are realized, where the agent’s payoff is $(\theta-p^{\prime})\mathbbm{1}\{a\neq a_{0}\}$, while the environment’s payoff is $p^{\prime}\mathbbm{1}\{a\neq a_{0}\}$.
Under ESEP, we have $\widetilde{\theta}=\theta$. Moreover, the agent’s action in the last period is determined as follows:
$a=\begin{cases}a^{*}&\left(p\leq\theta,p^{\prime}=p\right)\\ \overline{a}&\left(p>\theta,p^{\prime}=\varepsilon_{p}\right)\\ a_{0}&\left((p\leq\theta,p^{\prime}\neq p)\text{ OR }(p>\theta,p^{\prime}>\varepsilon_{p})\right)\end{cases}$
Note that the agent commits to such strategy in period 3 and 5.
Next, let’s consider the best response of the adaptive environment. Suppose that the principal offers $p$ in period 2 and she is now choosing $p^{\prime}$ in period 4. If $p\leq\theta$, then it is optimal to choose $p^{\prime}=p$, otherwise the agent would choose $a_{0}$. If $p>\theta$, then it is optimal for the principal to choose $p^{\prime}=\varepsilon_{p}$. Given this optimal behavior in Period 4, the payoff of the environment when offering $p$ in Period 2 is:
$p\Pr_{\pi}\left(\theta\geq p\right)+\varepsilon_{p}\Pr_{\pi}\left(\theta<p\right).$
Note that $\varepsilon_{p}$ is sufficiently small, with sufficiently large $T$, it is optimal for the principal to choose the monopoly price in Period 2. Moreover, whatever the value of $\varepsilon_{p}\in(0,\theta)$ is, the trade always happens in Period 5 (i.e., $a\neq a_{0}$ is chosen) in the equilibrium.
Lastly, if the environment is stationary, we have $p^{\prime}=p$. Then the trade happens if and only if $\theta\geq p$, and if it happens, the agent chooses the best action $a^{*}$. Therefore, ESEP has no-WER.
Theorem 3.
ESEP has no-WER, is partially safe, and welfare efficient.
Proof.
See Appendix B.7. ∎
Furthermore, we can show that ESEP with $\varepsilon_{p}\approx 0$ maximizes the “consumer surplus” among the class of all no-WER, partially safe, and welfare-efficient algorithms. See Appendix B.8 for details.
4 Discussion
Modeling Adversaries
Previous studies, such as Arora et al. (2012, 2018), have highlighted conceptual issues with the no-external-regret (no-ER) condition in non-stationary environments. In standard regret analysis, the benchmark payoff is computed under a fixed sequence of environment actions, making it difficult to interpret as the outcome of a counterfactual game where the environment also adapts to the learner’s actions. To address this issue, Arora et al. (2012) propose policy regret, a conceptually sound alternative. They show that meaningful regret bounds are unattainable without imposing constraints on the environment’s strategy space and derive regret bounds for $m$-memory-bounded adversaries.
In contrast, our approach restricts the environment’s strategy space by assuming that the environment optimizes its actions as a best response to the learner’s strategy, based on an economically relevant payoff function.
In what sense no-ER desirable?
The no-external-regret (no-ER) condition is a widely adopted criterion in modern machine learning, yet its desirability remains unclear. A common justification for its use is that no-ER algorithms also satisfy the weaker no-weak-external-regret (no-WER) condition, ensuring optimal learning in stationary environments. However, this justification assumes a non-strategic or stationary setting, an assumption that does not always hold in economic environments characterized by strategic interactions.
Without assuming a stationary environment, the connection between external regret and the time-averaged payoff becomes ambiguous. The results of this paper suggest that the no-ER condition imposes unnecessarily strong requirements on the algorithm compared to the no-WER condition, making it vulnerable to manipulation by adaptive and strategic opponents. Intuitively, the no-ER condition requires the algorithm to best respond whenever the opponent repeatedly plays the same action over a period of time. This allows the opponent to systematically probe the learner’s best responses and extract the full surplus.
These observations underscore the need for alternative algorithmic approaches that, while not satisfying the no-ER condition, are robust to strategic manipulation. A key challenge is to design learning algorithms that maintain strong performance in well-behaved environments while safeguarding against exploitation in adversarial settings.
Timing of moves in stage games
The current model assumes that the agent’s period-$t$ action distribution does not depend on the environment’s period $t$ action. This assumption is particularly relevant in settings where the agent interacts with a complex mechanism whose payment rule remains unclear even after reading its description—a reasonable first-order approximation of real-world online advertising auctions.
Practicality of ESEP
Although ESEP has welfare efficiency in addition to other two desirable properties (partial safety and no-WER) compared to EEP, it is worth noting that, for ESEP to be welfare efficient, it is necessary that the environment understands the structure of ESEP very well and responds optimally. In particular, the learner and the environment need to have a common understanding of the encoding of agent types. If the environment is not strategically sophisticated, ESEP is still no-WER and partially safe, but is not welfare-efficient anymore and incurs some additional losses for sublinear periods compared to EEP.
5 Concluding remarks
This paper examines the challenge of designing robust online learning algorithms for learners operating in potentially adaptive environments, particularly when learners possess private information. We demonstrate that while the widely used no-external-regret (no-ER) condition ensures strong learning performance in stationary environments, it renders algorithms highly vulnerable to strategic manipulation. In particular, we show that an adaptive opponent can systematically extract full surplus by eliciting the learner’s private information. Even learning algorithms satisfying a more refined regret notion, such as no-swap-regret, remain susceptible to this issue. Motivated by these vulnerabilities, we advocate shifting the focus toward partial safety, a design criterion aimed at preventing full surplus extraction by strategic opponents rather than minimizing regret against all possible sequences of opponent actions.
To this end, we introduce the Explore-Exploit-Punish (EEP) algorithm and establish that it achieves no-weak-external-regret (no-WER), ensuring effective learning in stationary environments while simultaneously providing partial safety in adaptive settings. This design philosophy embodies a balance between optimizing for the best-case scenario in well-behaved environments while guarding against worst-case outcomes in adversarial ones. We further examine the welfare implications of learning algorithms and propose Explore-Signal-Exploit-Punish (ESEP) as a welfare-efficient extension, highlighting the interplay between welfare efficiency and strategic sophistication.
Our findings have important implications for the design and application of online learning algorithms in economic contexts. They suggest that applying standard online learning algorithms without accounting for strategic interactions can be detrimental, potentially leading to exploitation and reduced payoffs for the learner. By prioritizing partial safety alongside no-WER, we provide a more robust framework for agents navigating uncertain and potentially adversarial environments.
Several promising directions for future research emerge from this work. A natural next step is to extend our analysis to multi-agent settings where multiple learners interact. Examining the robustness of our proposed algorithms in broader classes of games, including those with different payoff structures and/or information structures, would further enhance their applicability. Additionally, deriving lower bounds on the learning rate for algorithms that satisfy both partial safety and no-WER remains an open challenge. Intuitively, a tradeoff appears to exist between efficient learning and safety: to learn efficiently, the agent must adapt his actions over time based on his type. However, doing so may reveal his type in the early stages, potentially compromising the algorithm’s safety against an adaptive environment—as is indeed the case for Successive Elimination and UCB.
Another important avenue for exploration is the trade-off between safety and the range of environments in which an algorithm can effectively learn optimal actions. While we establish that partial safety is incompatible with no-ER but feasible with no-WER, a key open question is whether an algorithm can achieve partial safety while maintaining optimal learning performance across a broader class of environments beyond stationary settings. Addressing this question would provide deeper insights into the fundamental limitations and possibilities of robust learning in strategic environments.
Appendix
Appendix A Preliminaries
A.1 Doubling trick
Suppose that we have an online learning algorithm $\mathcal{L}$ with $(C,\delta,\gamma)$-no-ER (or no-WER). The doubling trick is a procedure to make an anytime algorithm $\overline{\mathcal{L}}$ by running $\mathcal{L}(\theta,T)$ repeatedly varying the choice of $T$ until the game ends, keeping the convergence rate $\delta$ and $\gamma$ the same. In particular, we can always make anytime $\gamma$-no-ER (or $\gamma$-no-WER) algorithm given a non-anytime anytime $\gamma$-no-ER (or $\gamma$-no-WER) algorithm.
1
Input: An online learning algorithm $\mathcal{L}(\theta,T)$
2 Initialize $T_{1}=1$;
3 Set $k=1$;
4
5while true do
6       Run algorithm $\mathcal{L}(\theta,T_{k})$ for $T_{k}$ rounds;
7       Double the horizon: $T_{k+1}=2T_{k}$;
8       Increment $k$: $k\leftarrow k+1$;
9
Algorithm 1 Doubling Trick
Lemma A.1 (Doubling trick).
Denote by $\overline{\mathcal{L}}$ the online learning algorithm obtained by applying the doubling trick to an online learning algorithm $\mathcal{L}$. Given $C>0$ , $\delta\in[0,1)$, and $\gamma\in(0,1)$, let
$C^{\prime}\coloneqq 2^{\delta}\frac{2^{\gamma+1}}{2^{\gamma}-1}C.$
  1. 1.
    If $\mathcal{L}$ has $(C,\delta,\gamma)$-no-WER, then $\overline{\mathcal{L}}$ is anytime $(C^{\prime},\delta,\gamma)$-no-WER.
  2. 2.
    If $\mathcal{L}$ has $(C,\delta,\gamma)$-no-ER and
    $\forall\theta\in\Theta,\ \forall T^{\prime}\leq T,\quad\mathrm{ER}(\mathcal{L}(\theta,T);\theta,T^{\prime},p_{1:T})\leq C\left(\log T\right)^{\delta}T^{\gamma},$ (7)
    then $\overline{\mathcal{L}}$ is anytime $(C^{\prime},\delta,\gamma)$-no-WER.
Proof.
Although this result is well-known in the literature (see Besson and Kaufmann (2018), for example), we provide a simple proof specific to our settings in Appendix B.2 for the sake of completeness. ∎
The condition (7) is imposed to guarantee the nonnegative external regret in the last epoch, as the external regret can be strictly negative in general. Popular no-ER algorithms such as EXP3 (Example 2) satisfy (7). A similar condition is automatically satisfied for weak external regret.161616 For any $\theta$, $t$, $\sigma_{\mathcal{A}}^{T}\in\mathcal{S}_{\mathcal{A}}^{T}$, and $\sigma_{P}\in\overline{\mathcal{S}}_{P}$, we have $\max_{a}\mathbbm{E}_{\sigma_{P}}\left[u(a,p_{t},\theta)\right]-\mathbbm{E}_{\sigma_{\mathcal{A}}^{T},\sigma_{P}}\left[u(a_{t},p_{t},\theta)\right]\geq 0.$
A.2 Confidence intervals
Lemma A.2 (Hoeffding’s inequality).
For any independent random variables $\mu_{1},\mu_{2},\ldots,\mu_{n}$ such that $a_{t}\leq\mu_{t}\leq b_{t}$ almost surely, and for their empirical mean $\overline{\mu}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mu_{t}$, the following inequality holds for all $\varepsilon>0$:
$\Pr\left(\overline{\mu}_{T}-\mathbb{E}[\overline{\mu}_{T}]\geq\varepsilon\right)\leq\exp\left(-\frac{2T^{2}\varepsilon^{2}}{\sum_{t=1}^{T}(b_{t}-a_{t})^{2}}\right).$
Similarly,
$\Pr\left(\overline{\mu}_{T}-\mathbb{E}[\overline{\mu}_{T}]\leq-\varepsilon\right)\leq\exp\left(-\frac{2T^{2}\varepsilon^{2}}{\sum_{t=1}^{T}(b_{t}-a_{t})^{2}}\right).$
Combining these, we get:
$\Pr\left(\left|\overline{\mu}_{T}-\mathbb{E}[\overline{\mu}_{T}]\right|\geq\varepsilon\right)\leq 2\exp\left(-\frac{2T^{2}\varepsilon^{2}}{\sum_{t=1}^{T}(b_{t}-a_{t})^{2}}\right).$ (8)
Corollary A.1.
Fix any $T\in\mathbb{Z}_{>0}$. Suppose that $(\mu_{t}(a))_{t=1}^{T}$ is iid-drawn from some $P(a)$ with mean $\mu(a)$, and $\mu_{t}(a)\in[0,1]$ for any $t\in T$ and $a\in\mathcal{A}$ almost surely. Let
$r_{t}(a)\coloneqq\sqrt{\frac{2\log T}{t}}.$
Then, for any $t\leq T$, we have
$\Pr_{P}\left(|\overline{\mu}_{t}(a)-\mu(a)|\geq r_{t}(a)\right)<\frac{2}{T^{4}}.$
Moreover, for any strategy $\sigma_{\mathcal{A}}$, $t$, and $a$, we have
$\Pr_{P,\sigma_{\mathcal{A}}}\left(|\overline{\mu}_{t}(a)-\mu(a)|\geq\rho_{t}(a)\right)<\frac{2}{T^{4}},$
where
$n_{t}(a)\coloneqq\sum_{s=1}^{t}\mathbbm{1}\{a_{t}=a\},\quad\rho_{t}(a)\coloneqq\sqrt{\frac{2\log T}{n_{t}(a)}}.$
Proof.
Fix any $t\leq T$. Let $a_{t}\equiv 0$, $b_{t}\equiv 1$, and $\varepsilon\coloneqq\sqrt{(2\log T)/t}$. Then the RHS of (8) becomes
$2\exp\left(-4\frac{T}{t}\log T\right)\leq 2\exp\left(-4\log T\right)=\frac{2}{T^{4}}.$
The confidence intervals are defined as follows (see also Example 4):
$\begin{array}[]{rlrl}n_{t}(a)&\coloneqq\sum_{s=1}^{t}\mathbbm{1}\{a_{s}=a\},&\hat{\mu}_{t}(a,\theta)&\coloneqq\frac{1}{n_{t}(a)}\sum_{s=1}^{t}\mathbbm{1}\{a_{s}=a\}u(a,p_{s},\theta),\\ \rho_{t}(a)&\coloneqq\sqrt{\frac{2\log T}{n_{t}(a)}},&&\\ \ell_{t}(a,\theta)&\coloneqq\hat{\mu}_{t}(a,\theta)-\rho_{t}(a),&u_{t}(a,\theta)&\coloneqq\hat{\mu}_{t}(a,\theta)+\rho_{t}(a).\end{array}$ (12)
Corollary A.2.
Fix any $T\in\mathbb{Z}_{>0}$ such that $T\geq|\mathcal{A}|$. Suppose that $(\mu_{t}(a))_{t=1}^{T}$ is iid-drawn from some $P(a)$ with mean $\mu(a)$, and $\mu_{t}(a)\in[0,1]$ for any $t\in T$ and $a\in\mathcal{A}$ almost surely. Then, for any strategy $\sigma_{\mathcal{A}}$, we have
$\Pr_{P,\sigma_{\mathcal{A}}}\bigl{(}\forall a\in\mathcal{A}\ \forall t\leq T,\quad|\mu_{t}(a)-\mu(a)|\leq\rho_{t}(a)\bigr{)}\geq 1-\frac{2}{T^{2}}.$ (13)
Proof.
$\Pr_{P,\sigma_{\mathcal{A}}}\bigl{(}\forall a\in\mathcal{A}\ \forall t\leq T,\quad|\mu_{t}(a)-\mu(a)|\leq\rho_{t}(a)\bigr{)}$ $=1-\Pr_{P,\sigma_{\mathcal{A}}}\bigl{(}\exists a\in\mathcal{A}\ \exists t\leq T,\quad|\mu_{t}(a)-\mu(a)|>\rho_{t}(a)\bigr{)}$ (14)
$=1-\Pr_{P,\sigma_{\mathcal{A}}}\left(\bigcup_{a\in\mathcal{A}}\bigcup_{t\leq T}\left\{|\mu_{t}(a)-\mu(a)|>\rho_{t}(a)\right\}\right)$ (15)
$\geq 1-\sum_{a\in\mathcal{A}}\sum_{t\leq T}\Pr\bigl{(}|\mu_{t}(a)-\mu(a)|>\rho_{t}(a)\bigr{)}$ (16)
$\geq 1-|\mathcal{A}|T\frac{2}{T^{4}}$ (17)
$\geq 1-\frac{2}{T^{2}}\quad(\because\ |\mathcal{A}|\leq T)$ (18)
A.3 No-WER algorithms
Throughout this section, we fix the agent’s type $\theta\in\Theta$ and subsume it. We also assume $x_{t}(a)\equiv 1$ for all $t$ and $a\neq a_{0}$. Denote $u(a,(1,p),\theta)$ by $u(a,p)$. Assume that the environment’s actions $(p_{t})_{t=1}^{T}$ is iid-drawn from $\sigma_{P}\in\mathcal{M}$. Let $r_{t}(a)\coloneqq u(a,p_{t})+1/2\in[0,1]$ (Here, $r$ stands for “reward”.) Then, $(r_{t}(a))_{t}$ is iid-drawn from $P_{a}\in\Delta([0,1])$ with mean $\mu(a)\in[0,1]$ for each $a\in\mathcal{A}$. Let
$a^{*}\in\mathop{\rm arg~{}max}\limits_{a\in\mathcal{A}}\mu(a),\quad\Delta(a)\coloneqq\mu(a^{*})-\mu(a)\geq 0.$
Note that for any strategy $\sigma_{\mathcal{A}}$, we have
$\mathrm{WER}(\sigma_{\mathcal{A}};T,\sigma_{P})=\mathbbm{E}_{\sigma_{\mathcal{A}}}\left[\sum_{t=1}^{T}\Delta(a_{t})\right]=\sum_{a\in\mathcal{A}}\Delta(a)n_{T}(a),$
where $n_{T}(a)$ is defined in (12). Note that $n_{T}(a)$ is a random variable depending on $\sigma_{\mathcal{A}}$.
Definition A.1 (Clean Events).
Suppose that the confidence intervals are defined as (12). For a fixed strategy $\sigma_{\mathcal{A}}\in\mathcal{S}_{\mathcal{A}}$ and stationary strategy $\sigma_{P}\in\mathcal{S}_{\mathcal{M}}^{0}$, the following event is called the clean event (CE):
$\left\{\forall a\in\mathcal{A}\ \forall t\leq T,\quad|\mu_{t}(a)-\mu(a)|\leq\rho_{t}(a)\right\}.$ (19)
Conditional on CE, for each action, the confidence interval includes the true mean throughout the game play.
Remark A.1.
Corollary A.2 states that the clean event happens with probability at least $1-2/T^{2}$.
Lemma A.3.
Fix any strategy $\sigma_{\mathcal{A}}\in\mathcal{S}_{\mathcal{A}}$, horizon $T\in\mathbb{Z}_{>0}$, and stationary strategy $\sigma_{P}\in\mathcal{S}_{\mathcal{M}}^{0}$. Suppose that, conditional on the clean event, we have
$\forall a\in\mathcal{A},\quad\Delta(a)\leq O\left(\sqrt{\frac{\log T}{n_{T}(a)}}\right).$ (20)
Then, we have $\mathrm{WER}(\sigma_{\mathcal{A}};T,\sigma_{P})\leq O(\sqrt{|\mathcal{A}|T\log T})$.
Proof.
We will subsume $\sigma_{\mathcal{A}}$ and $\sigma_{P}$ throughout the proof. By (20), we have
$\mathbbm{E}\left[\mathrm{WER}(T)\mid\mathrm{CE}\right]$ $=\sum_{a}\Delta(a)n_{T}(a)\leq O(\sqrt{\log T})\sum_{a}\sqrt{n_{T}(a)}.$ (21)
By Jensen’s inequality, we have
$\sum_{a}\sqrt{n_{T}(a)}=|\mathcal{A}|\sum_{a}\frac{1}{|\mathcal{A}|}\sqrt{n_{T}(a)}\leq|\mathcal{A}|\sqrt{\sum_{a}\frac{n_{T}(a)}{|\mathcal{A}|}}=\sqrt{|\mathcal{A}|T}.$
Therefore, we have
$\mathrm{WER}(T)$ $=\mathbbm{E}[\mathrm{WER}(T)\mid\mathrm{CE}]\Pr(\mathrm{CE})+\mathbbm{E}[\mathrm{WER}(T)\mid\lnot\mathrm{CE}]\Pr(\lnot\mathrm{CE})$ (22)
$\leq O\left(\sqrt{|\mathcal{A}|T\log T}\right)+\frac{1}{T^{2}}T$ (23)
$=O\left(\sqrt{|\mathcal{A}|T\log T}\right).$ (24)
Proposition A.1.
Successive Elimination (Algorithm 3) has $\left(5\sqrt{|\mathcal{A}|},\frac{1}{2},\frac{1}{2}\right)$-no-WER.
Proof.
First, suppose that the clean event happens. Suppose that action $a$ is eliminated at the end of period $t$. Conditional on the clean event (CE), we have
$\Delta(a)$ $\leq 2(\rho_{t}(a)+\rho_{t}(a^{*}))=4\rho_{t}(a)=4\sqrt{\frac{\log T}{n_{t}(a)}}=4\sqrt{\frac{\log T}{n_{T}(a)}},$ (25)
where the last equality follows since $n_{t}(a)=n_{T}(a)$ if $a$ is removed at the end of period $t$. Then, we have
$\mathrm{WER}(T)$ $=\sum_{a}\Delta(a)n_{T}(a)$ (26)
$=\mathbbm{E}\left[\sum_{a}\Delta(a)n_{T}(a)\mid\mathrm{CE}\right]\Pr(\mathrm{CE})+\mathbbm{E}\left[\sum_{a}\Delta(a)n_{T}(a)\mid\lnot\mathrm{CE}\right]\Pr(\lnot\mathrm{CE})$ (27)
$\leq 4\sqrt{\log T}\sum_{a}\sqrt{n_{T}(a)}+\frac{1}{T^{2}}T$ (28)
$\leq 4\sqrt{\log T}\sqrt{|\mathcal{A}|T}+\frac{1}{T}$ (29)
$=5\sqrt{|\mathcal{A}|\log T}\sqrt{T}.$ (30)
Proposition A.2.
UCB has $\frac{1}{2}$-no-WER.
Proof.
Denote the arm pulled in period $t$ by $a_{t}$. Condition on the CE. For any $t\in[T]$ and $a_{t}\in A$, by the construction of CI, we have
$\mu(a_{t})+\rho_{t}(a_{t})$ $\geq\hat{\mu}_{t}(a_{t}),\quad u_{t}(a^{*})\geq\mu(a^{*}).$ (31)
Then,
$\mu(a_{t})+2\rho_{t}(a_{t})$ $\geq\mu_{t}(a_{t})+\rho_{t}(a_{t})$ (32)
$=u_{t}(a_{t})$ (33)
$=u_{t}(a^{*})\quad(\because\ \text{UCB})$ (34)
$\geq\mu(a^{*}).$ (35)
Thus, we have
$\Delta(a_{t})\leq 2\rho_{t}(a_{t})=O\left(\sqrt{\frac{\log T}{n_{t}(a_{t})}}\right).$ (36)
For any $a\neq a^{*}$, let $t$ be the last period in which arm $a$ is pulled, so we have $n_{t}(a)=n_{T}(a)$. Then, (20) holds for any $a\neq a^{*}$. By Lemma A.3, we have the result. ∎
There is another more primitive no-WER algorithm called Uniform Exploration: try each arm $N$ times, and then choose the arm with the highest average reward in all remaining rounds. If we choose $N$ suitably to balance the exploration-exploitation trade-off, it achieves no-weak-ER.
Proposition A.3.
If we set $N\coloneqq T^{2/3}(\log T)^{1/3}$, then Uniform Exploration has $\frac{2}{3}$-no-WER.
Proof.
Let $\overline{\rho}_{N}\coloneqq\sqrt{(2\log T)/N}$ and $K\coloneqq|\mathcal{A}|$.
For each action, true mean $\mu(a)$ is included in the confidence interval $\overline{\mu}_{N}(a)\pm\overline{\rho}_{N}$ with probability at least $1-2/T^{4}$. Thus, the clean event (CE), on which the true means are in the confidence intervals for all actions, happens with probability $1-O(T^{4})$. Conditional on the CE, if action $a$ is chosen at the end of the exploration phase, we have $\mu(a^{*})-\mu(a)\leq 2\overline{\rho}_{N}$. Therefore, conditional on CE, we have the regret bound
$KN+(T-KN)2\overline{\rho}_{N}.$
Therefore, by setting $N\coloneqq T^{2/3}(\log T)^{1/3}$, we have the following regret bound
$R(T)$ $\leq\left(1-O(T^{4})\right)\left[KN+(T-KN)2\overline{\rho}_{N}\right]+O(T^{4})T$ (37)
$=O\left(T^{2/3}\left(\log T\right)^{1/3}\right).$ (38)
A.4 Examples of No-WER algorithms
1
Input: Horizon $T$, agent’s type $\theta$, action set $\mathcal{A}$, learning rate $\eta_{t}>0$
2 Rescale $u$ so that its range is $[0,1]$;
3 Initialize weights $w_{1}(a)\coloneqq 1/|\mathcal{A}|$ for each $a\in\mathcal{A}$
4for $t\in\{1,\dots,T\}$ do
5       Let ${q_{t}(a)=\frac{w_{t}(a)}{\sum_{a}w_{t}(a)}}\text{ for }a\in\mathcal{A}$;
6       Draw action $a_{t}$ from the multinomial distribution $\left(q_{t}(a)\right)_{a\in\mathcal{A}}$;
7       Observe $(x_{t}(a_{t}),p_{t}(a_{t}))$. Let $u_{t}\coloneqq u(a_{t},(x_{t},p_{t}),\theta)$;
8       for $a\in\mathcal{A}$ do
9             if $a=a_{t}$ then
10                  $w_{t+1}(a)=w_{t}(a)\exp\left(\eta_{t}\cdot\left(1-\frac{1}{q_{t}(a)}(1-u_{t})\right)\right)$
11            if $a\neq a_{t}$ then
12                  $w_{t+1}(a)=w_{t}(a)$
13
14
15
Algorithm 2 EXP3
Input: Horizon $T$, agent’s type $\theta$, action set $\mathcal{A}$
1 Rescale $u$ so that its range is $[0,1]$;
2 Initialize the active action set $\mathcal{A}_{\mathrm{active}}\coloneqq\mathcal{A}$ and time counter $t\coloneqq 1$;
3
4while $t\leq T$ do
5       for $a\in\mathcal{A}_{\mathrm{active}}$ do
6             Choose arm $a_{t}=a$;
7             Observe $u_{t}\coloneqq u(a_{t},(x_{t}(a_{t}),p_{t}(a_{t})),\theta)$; Compute $n_{t}(a)$, $\hat{\mu}_{t}(a)$, $\mathrm{LCB}_{t}(a)$, and $\mathrm{UCB}_{t}(a)$ ;
8             $t\leftarrow t+1$;
9
10
11      for $a\in\mathcal{A}_{\mathrm{active}}$ do
12             if $\exists a^{\prime}\in\mathcal{A}_{\mathrm{active}},\ \mathrm{LCB}_{t}(a^{\prime})>\mathrm{UCB}_{t}(a)$ then
13                   $\mathcal{A}_{\mathrm{active}}\leftarrow\mathcal{A}_{\mathrm{active}}\setminus\{a\}$;
14
15
16
17
Algorithm 3 Successive Elimination
Input: Horizon $T$, agent’s type $\theta$, action set $\mathcal{A}$
1 Rescale $u$ so that its range is $[0,1]$;
2 Initialize $t\coloneqq 1$;
3
// Cold start
4 for $a\in\mathcal{A}_{\mathrm{active}}$ do
5       Choose arm $a_{t}=a$;
6       Observe $u_{t}\coloneqq u(a_{t},(x_{t},p_{t}),\theta)$; Compute $n_{t}(a)$, $\hat{\mu}_{t}(a)$, $\mathrm{LCB}_{t}(a)$, and $\mathrm{UCB}_{t}(a)$;
7       $t\leftarrow t+1$;
8
9
// Main loop
10 while $t\leq T$ do
11       Choose arm $a_{t}=\arg\max_{a\in\mathcal{A}}\mathrm{UCB}_{t-1}(a)$;
12       Observe $u(a_{t},(x_{t},p_{t}),\theta)$; Compute $n_{t}(a)$, $\hat{\mu}_{t}(a)$, $\mathrm{LCB}_{t}(a)$, and $\mathrm{UCB}_{t}(a)$;
13       $t\leftarrow t+1$;
14
15
Algorithm 4 UCB
Appendix B Omitted proofs
B.1 Proof of Proposition 1
It is well known that these three algorithms are no-WER (see Appendix A.3.)
For UE and SE, consider the following strategy of the adaptive environment: for some $a_{1}\neq a_{0}$, for any period $t\leq T_{1}$ before the learner drops all actions except $a_{1}$ at the end of period $T_{1}$,
$(x_{t}(a),p_{t}(a))=\begin{cases}(1,1/2)&(a=a_{1})\\ (0,0)&(o.w.)\end{cases}.$
Both UE and SE remove all actions but $a_{1}$ with sublinear $T_{1}$. For $t\geq T_{1}+1$, the environment sets $p_{t}(a_{1})=1/2$, so she achieves $1/2+o(1)$ average payoff. This proves that UE and SE are unsafe.
For UCB, suppose that there are $K$ possible agent types $\{\theta_{1},\dots,\theta_{K}\}$ with $20\leq\theta_{1}<\theta_{2}<\dots<\theta_{K}<1/$ 2. By assumption, we have $|\mathcal{A}|\geq K+1$. Let $\varepsilon\coloneqq\max_{k\in[K]}|\theta_{k+1}-\theta_{k}|>0$, where $\theta_{K+1}\coloneqq 1/2$.
For sufficiently large $T$, for the early periods, UCB chooses least explored actions (i.e., chooses action $a\in\arg\min_{a\in\mathcal{A}}n_{t}(a)$ in period $t$) as the effect of the confidence intervals dominates. If the number of explorations is the same for all actions, then UCB chooses the one with the best past performance.
Let $M\coloneqq|\mathcal{A}|$ and $K\coloneqq|\Theta|$. Note that $M\geq K+1$ by assumption. Consider the following environment’s strategy: for $t\leq M$,
$(x_{t}(a_{k}),p_{t}(a_{k}))=\begin{cases}(1,\theta_{k}-\varepsilon)&(k\in\{1,\dots,K\})\\ (0,0)&(k\in\{K+1,\dots,M\}),\end{cases}$
and for $t\in\{M+1,\dots,2M\}$, $x_{t}(a)=p_{t}(a)=0$ for all $a\in\mathcal{A}$.
For $t\in\{1,\dots,M\}$, UCB chooses each of the $M$ actions once. Then, for $t\in\{M+1,\dots,2M\}$, UCB chooses actions in descending order of past performance observed during $t\leq M$. Define $a_{0},a_{K+1},\dots a_{M}$ as null actions. A learner with type $\theta_{k}$ selects actions in the order $a_{1},a_{2},\dots,a_{k}$, followed by the null actions, and then $a_{k+1},\dots,a_{M}$. Therefore, by observing the behavior during periods $t\in\{M+1,\dots,2M\}$, the environment can perfectly learn the agent’s type $\theta$. For the remainder of the game, the environment sets $x_{t}(a_{k})=1,p_{t}(a_{k})=\theta-\varepsilon_{0}$ with small enough $\varepsilon_{0}$ for all $k\geq 1$.
B.2 Proof of Lemma A.1
Fix any $\theta$ and $T$. Let
$S_{m}\coloneqq\sum_{k=0}^{m}2^{k}=2^{m+1}-1,\quad T_{m}\coloneqq\sum_{k=0}^{m}\left(\log(2^{k})\right)^{\delta}(2^{k})^{\gamma}.$
Let $M$ be the integer such that $S_{M-1}\leq T<S_{M}$. By condition (7), the external regret of $\overline{L}(\theta)$ is bounded from above by $CT_{M}$.
$T_{M}$ $=\sum_{k=0}^{M}\left(\log(2^{k})\right)^{\delta}(2^{k})^{\gamma}=\left(\log 2\right)^{\delta}\sum_{k=0}^{M}k^{\delta}(2^{\gamma})^{k}$ (39)
$\leq M^{\delta}\left(\log 2\right)^{\delta}\sum_{k=0}^{M}(2^{\gamma})^{k}$ (40)
$=M^{\delta}\left(\log 2\right)^{\delta}\frac{1}{2^{\gamma}-1}\left(2^{\gamma(M+1)}-1\right).$ (41)
Since $S_{M-1}\leq T$, we have $M\log 2\leq\log(T+1)$, and thus
$1\displaystyle 2^{\gamma(M+1)}-$ 1 $=2^{\gamma}\exp\left(\gamma M\log 2\right)-1=2^{\gamma}\exp\left(\gamma\log(T+1)\right)-1=2^{\gamma}(T+1)^{\gamma}-1$ (42)
$<2^{\gamma}T^{\gamma}+(2^{\gamma}-1)$ (43)
$\leq 2^{\gamma+1}T^{\gamma},$ (44)
where the second last inequality holds since $(T+1)^{\gamma}<T^{\gamma}+1$ for any $T>0$ and $\gamma\in(0,1)$. We also have
$M^{\delta}\left(\log 2\right)^{\delta}\leq\left(\log(T+1)\right)^{\delta}\leq\left(1+\log T\right)^{\delta}<\left(2\log T\right)^{\delta},$
where the last inequality holds since we assume $T\geq 3$. Therefore, we have
$T_{M}<2^{\delta}\frac{2^{\gamma+1}}{2^{\gamma}-1}\left(\log T\right)^{\delta}T^{\gamma}.$
B.3 Proof of Lemma 1
Assume that an online learning algorithm $\mathcal{L}$ has no-ER with regret upper bound $R$. Fix any $a$, $T$, $\theta$, and $\sigma_{P}\in\overline{\mathcal{S}}_{P}$. Let $\sigma_{A}^{T}\coloneqq\mathcal{L}(\theta,T)$. We have
$\mathop{\mathbbm{E}}_{\sigma_{P}}\left[\sum_{t=1}^{T}u(a,(x_{t},p_{t}),\theta)\right]-\mathop{\mathbbm{E}}_{\sigma_{A}^{T},\sigma_{P}}\left[\sum_{t=1}^{T}u(a_{t},(x_{t},p_{t}),\theta)\right]$ (45)
$=\mathbbm{E}\left[\mathbbm{E}\left[\sum_{t=1}^{T}u(a,(x_{t},p_{t}),\theta)-\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T}u(a_{t},(x_{t},p_{t}),\theta)\right]\mid(x_{1:T},p_{1:T})\right]\right].$ (46)
Since $\mathcal{L}$ has no-ER, for any $x_{1:T},p_{1:T}$, we have
$\sum_{t=1}^{T}u(a,(x_{t},p_{t}),\theta)-\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T}u(a_{t},(x_{t},p_{t}),\theta)\right]\leq R(T).$
Therefore, the RHS of (46) is bounded from above by $R(T)$, with $R(T)=o(T)$ by assumption. ∎
B.4 Proof of Theorem 1
Let $\Theta\coloneqq\{\theta_{1},\dots,\theta_{K}\}$, where $20\leq\theta_{1}<\theta_{2}<\cdots<\theta_{K}<1/$ 2. In this section, we show the following theorem, which implies Theorem 1.
Theorem A.1.
Suppose that the agent uses a no-ER algorithm and agent’s type is $\theta_{k}\in\Theta$. Then, for any $\varepsilon>0$, there exists a principal’s strategy such that the principal obtains the average ex-ante expected payoff at least $(1-\varepsilon)\theta_{k}+o(1)$.
Notation
For $A\subseteq\mathcal{A}$ and $T^{\prime}\leq T$, let $n_{T^{\prime}}(A)\coloneqq\mathbbm{E}\left[\sum_{t=1}^{T^{\prime}}\mathbbm{1}\{a_{t}\in A\}\right]$, where the expectation is taken given a probability distribution over $(a_{1:T},x_{1:T},p_{1:T})$. Define $n_{T^{\prime}}(a)\coloneqq n_{T^{\prime}}(\{a\})$ and $n_{T^{\prime}}(\lnot a)\coloneqq n_{T^{\prime}}(A\setminus\{a\})$. Let $\delta_{k}\coloneqq(\theta_{k}-\theta_{k-1})/2>0$ for $k\in[K]$.
B.4.1 Proof of Theorem A.1
Suppose that $\mathcal{L}$ is a no-ER algorithm. Assume for simplicity that $T_{k}$ is odd for all $k\in[K]$. [^17] When $T_{k}$ is even, we need to decide how to break ties.
Lemma A.4.
Denote the agent’s private type by $\theta\in\Theta$. For any $\varepsilon^{\prime}>0$, there exists $T_{K}=o(T)$ such that for any $\theta\in\Theta$, the principal can learn whether $\theta\geq\theta_{K}$ or not with probability at least $1-\varepsilon^{\prime}$ at the end of period $T_{K}$ by the following strategy: let $T_{K}\coloneqq\frac{2R(T)}{\varepsilon^{\prime}\delta_{K}}$, and for $t\leq T_{K}$, let
$x_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a})\\ 1&(a=\overline{a})\end{cases},\quad p_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a})\\ \theta_{K}-\delta_{K}&(a=\overline{a})\end{cases}.$
If $n_{T_{K}}(\overline{a})\geq T_{K}/2$, then conclude $\theta=\theta_{K}$; otherwise, conclude $\theta\leq\theta_{K-1}$.
Proof.
Choose $\overline{a}\in\mathcal{A}\setminus\{a_{0}\}$. Fix any $T_{K}\leq T$. First, consider a principal’s strategy such that
$x_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a},\text{ or }t\geq T_{K}+1)\\ 1&(a=\overline{a},\text{ and }t\leq T_{K})\end{cases},\quad p_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a},\text{ or }t\geq T_{K}+1)\\ \theta_{K}-\delta_{K}&(a=\overline{a},\text{ and }t\leq T_{K})\end{cases}.$
Throughout the proof, we fix this $x_{1:T}$ and $p_{1:T}$.
Case (i): $\theta=\theta_{K}$.
By the no-ER condition, we have
$\sum_{t=1}^{T_{K}}u\left(\overline{a},(x_{t},p_{t}),\theta\right)-\mathbbm{E}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T_{K}}u\left(a_{t},(x_{t},p_{t}),\theta\right)\right]\leq R(T)=o(T).$ (47)
As for the LHS, we have
$\sum_{t=1}^{T_{K}}u\left(\overline{a},(x_{t},p_{t}),\theta\right)-\mathbbm{E}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T_{K}}u\left(a_{t},(x_{t},p_{t}),\theta\right)\right]$ (48)
$=\sum_{t=1}^{T_{K}}\left(\theta-(\theta_{K}-\delta_{K})\right)-\mathbbm{E}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T_{K}}\mathbbm{1}\{a_{t}=\overline{a}\}(\theta-(\theta_{K}-\delta_{K}))\right]$ (49)
$=\mathbbm{E}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T_{K}}\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-(\theta_{K}-\delta_{K}))\right]$ (50)
$=\sum_{t=1}^{T_{K}}\mathbbm{E}\left[\mathbbm{E}\left[\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-(\theta_{K}-\delta_{K}))\mid a_{1:t-1}\right]\right]$ (51)
$=\delta_{K}\mathbbm{E}\left[\mathbbm{E}\left[n_{T_{K}}(\lnot\overline{a})\mid a_{1:t-1}\right]\right]\quad(\because\ \theta=\theta_{K})$ (52)
$=\delta_{K}\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}[n_{T_{K}}(\lnot\overline{a})].$ (53)
Thus, we have
$\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}[n_{T_{K}}(\lnot\overline{a})]\leq\frac{R(T)}{\delta_{K}}.$ (54)
By Markov’s inequality, we have
$\Pr\left(n_{T_{K}}(\lnot\overline{a})<\frac{T_{K}}{2}\right)\geq 1-\frac{2}{T_{K}}\frac{R(T)}{\delta_{K}}.$ (55)
Let $T_{K}\coloneqq\frac{2R(T)}{\varepsilon^{\prime}\delta_{K}}=o(T)$. Then, we have
$\Pr\left(n_{T_{K}}(\lnot\overline{a})<\frac{T_{K}}{2}\right)\geq 1-\varepsilon^{\prime}.$
Case (ii): $\theta\leq\theta_{K-1}$
By the no-ER condition, we have
$-\mathbbm{E}_{\sigma_{A}^{T}}\left[\sum_{t=1}^{T_{K}}u\left(a_{t},(x_{t},p_{t}),\theta\right)\right]\leq R(T)=o(T).$ (56)
By the similar argument to Case (i), we have
$\delta_{K}\mathbbm{E}_{\sigma_{A}^{T}}\left[n_{T_{K}}(\overline{a})\right]\leq R(T),$
and thus
$\Pr\left(n_{T_{K}}(\overline{a})<\frac{T_{K}}{2}\right)\geq 1-\varepsilon^{\prime},$
with $T_{K}\coloneqq\frac{2R(T)}{\varepsilon^{\prime}\delta_{K}}=o(T)$. ∎
Lemma A.5.
Fix any $\varepsilon^{\prime}>0$. Let
$T_{K}\coloneqq\frac{2}{\varepsilon^{\prime}\delta_{K}}R(T),\quad T_{K-1}\coloneqq\frac{2}{\varepsilon^{\prime}(1-\varepsilon^{\prime})\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right).$
Consider the following strategy: for $t\in\{1,\dots,T_{K}+T_{K-1}\}$,
$x_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a})\\ 1&(a=\overline{a})\end{cases},\quad p_{t}(a)\coloneqq\begin{cases}0&(a\neq\overline{a})\\ \theta_{K}-\delta_{K}&(a=\overline{a},t\leq T_{K})\\ \theta_{K-1}-\delta_{K-1}&(a=\overline{a},T_{K}+1\leq t\leq T_{K}+T_{K-1})\end{cases}.$
Suppose that $\theta\leq\theta_{K-1}$. There exists $T_{K-1}=o(T)$ such that at the end of period $T_{K}+T_{K-1}$, with probability at least $(1-\varepsilon^{\prime})^{2}$, the principal can learn whether $\theta\geq\theta_{K-2}$ or not.
Proof.
Suppose that $\theta\leq\theta_{K-1}$. Define the $K$-clean event as
$\mathrm{CE}_{K}\coloneqq\{\text{Principal correctly learn $\theta\leq\theta_{K-1}$ at the end of period $T_{K}$}\},$
which happens with probability at least $1-\varepsilon^{\prime}$ by Lemma A.4. Let $I_{K}\coloneqq\{1,.\dots,T_{K}\}$, $I_{K-1}\coloneqq\{T_{K}+1,\dots,T_{K}+T_{K-1}\}$, and $n_{I}(A)\coloneqq\sum_{t\in I}\mathbbm{1}\{a_{t}\in A\}$ for any $I\subseteq[T]$.
Case (i): $\theta=\theta_{K-1}$
By a similar argument to the one in the proof of Lemma A.4, the no-ER condition for action $\overline{a}$ implies
$-\delta_{K}\mathop{\mathbbm{E}}_{\sigma_{A}^{T},\mathbbm{P}}\left[n_{I_{K}}(\lnot\overline{a})\right]+\delta_{K-1}\mathop{\mathbbm{E}}_{\sigma_{A}^{T},\mathbbm{P}}\left[n_{I_{K-1}}(\lnot\overline{a})\right]\leq R(T),$
and thus
$\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\right]$ $\leq\frac{1}{\delta_{K-1}}\left(R(T)+\delta_{K}\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K}}(\lnot\overline{a})\right]\right)$ (57)
$\leq\frac{1}{\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right).$ (58)
Regarding the LHS, we have
$\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\right]$ $=\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\mid\mathrm{CE}_{K}\right]\Pr(\mathrm{CE}_{K})+\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\mid\lnot\mathrm{CE}_{K}\right]\Pr(\lnot\mathrm{CE}_{K})$ (59)
$\geq\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\mid\mathrm{CE}_{K}\right](1-\varepsilon^{\prime}).$ (60)
We then have
$\mathop{\mathbbm{E}}_{\sigma_{A}^{T}}\left[n_{I_{K-1}}(\lnot\overline{a})\mid\mathrm{CE}_{K}\right]\leq\frac{1}{1-\varepsilon^{\prime}}\frac{1}{\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right).$ (61)
By Markov’s inequality, we have
$\Pr\left(n_{I_{K-1}}(\lnot\overline{a})\leq\frac{T_{K-1}}{2}\mid\mathrm{CE}_{K}\right)$ $\geq 1-\frac{2}{T_{K-1}}\frac{1}{1-\varepsilon^{\prime}}\frac{1}{\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right)\geq 1-\varepsilon^{\prime},$ (62)
with $T_{K-1}\coloneqq\frac{2}{\varepsilon^{\prime}(1-\varepsilon^{\prime})\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right)$.
Case (ii): $\theta\leq\theta_{K-2}$
By the no-ER condition for action $a_{0}$, we have
$\left(\theta_{K}-(\theta+\delta_{K})\right)\mathbbm{E}\left[n_{I_{K}}(\overline{a})\right]+\left(\theta_{K-1}-(\theta+\delta_{K-1})\right)\mathbbm{E}\left[n_{I_{K-1}}(\overline{a})\right]\leq R(T).$ (63)
This implies
$\mathbbm{E}\left[n_{I_{K-1}}(\overline{a})\right]\leq\frac{1}{\delta_{K-1}}R(T)\leq\frac{1}{\delta_{K-1}}(R(T)+\delta_{K}T_{K}).$ (64)
By the same argument as in Case (i), we have
$\Pr\left(n_{I_{K-1}}(\overline{a})\leq\frac{T_{K-1}}{2}\mid\mathrm{CE}_{K}\right)$ $\geq 1-\varepsilon^{\prime},$ (65)
with $T_{K-1}\coloneqq\frac{2}{\varepsilon^{\prime}(1-\varepsilon^{\prime})\delta_{K-1}}\left(R(T)+\delta_{K}T_{K}\right)$.
Corollary A.3.
Fix any $\varepsilon^{\prime}>0$. Let
$T_{k}\coloneqq\frac{2}{\varepsilon^{\prime}}\frac{1}{(1-\varepsilon^{\prime})^{K-k}}\left(R(T)+\sum_{s=k+1}^{K}\delta_{s}T_{s}\right).$
For any $\theta_{k}\in\Theta$, at the end of period $S_{k}\coloneqq\sum_{s=k}^{K}T_{s}$, with probability at least $(1-\varepsilon^{\prime})^{K-k+1}$, the principal learns $\theta=\theta_{k}$. Moreover, $S_{k}=o(T)$ for any $k\in[K]$.
Lemma A.6.
Fix any $T$, $\varepsilon^{\prime}$, and $(\Delta_{k})_{k}\in\mathbb{R}_{>0}^{K}$. Suppose that the true mean is $\theta_{k}\in\Theta$. If the principal uses Algorithm 5, then with probability at least $(1-\varepsilon^{\prime})^{K}$, her undiscounted payoff is $\theta_{k}-\Delta_{k}-o(1)$.
Theorem A.2.
Suppose that the agent uses a no-ER algorithm and the true agent’s type is $\theta_{k}\in\Theta$. For any $\varepsilon\in(0,1)$, Algorithm 5 with
$\varepsilon^{\prime}\coloneqq 1-\exp\left(\frac{\log\left(\varepsilon/2\right)}{K}\right)$
learns that $\theta=\theta_{k}$ at the end of period $T_{K}+\cdots+T_{k}=o(T)$ with probability at least $(1-\varepsilon)$. Moreover, with
$\Delta_{k}\coloneqq\frac{\varepsilon}{2-\varepsilon}\theta_{k},$
the principal can get the average payoff of $(1-\varepsilon)\theta_{k}+o(1)$.
Proof.
Note that $\Delta_{k}$ is chosen so that
$\left(1-\frac{\varepsilon}{2}\right)(\theta_{k}-\Delta_{k})\geq(1-\varepsilon)\theta_{k}.$
Let $S_{k}\coloneqq T_{K}+\cdots T_{k}$ and $J_{k}\coloneqq\{S_{k}+1,\dots,T\}$. Given $\theta_{k}\in\Theta$, we define the clean event $\mathrm{CE}$ as the event on which the principal learns $\theta=\theta_{k}$ correctly at the end of period $S_{k}$. To claim the second part of the statement, it suffices to show that
$\mathbbm{E}\left[n_{J_{k}}(\overline{a})\mid\mathrm{CE}\right]\geq T-o(T).$ (66)
Note that there exists $\overline{\theta}\in[0,1/2]$ such that, for any $k$, $\theta_{k}\leq\overline{\theta}$. By the same argument as before, the no-ER condition for action $\overline{a}$ implies
$\mathbbm{E}\left[\sum_{t\in J_{k}}\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-p_{t}(\overline{a}))\right]\leq R(T)+\overline{\theta}S_{k}=o(T).$ (67)
Regarding the LHS, we have
$\mathbbm{E}\left[\sum_{t\in J_{k}}\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-p_{t}(\overline{a}))\right]$ (68)
$=\mathbbm{E}\left[\sum_{t\in J_{k}}\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-p_{t}(\overline{a}))\mid\mathrm{CE}\right]\Pr(\mathrm{CE})+\mathbbm{E}\left[\sum_{t\in J_{k}}\mathbbm{1}\{a_{t}\neq\overline{a}\}(\theta-p_{t}(\overline{a}))\mid\lnot\mathrm{CE}\right]\Pr(\lnot\mathrm{CE})$ (69)
$\geq(1-\varepsilon^{\prime})^{K}\Delta_{k}\mathbbm{E}\left[n_{J_{k}}(\lnot\overline{a})\mid\mathrm{CE}\right],$ (70)
where the inequality follows since the second term of the second line is positive by construction. We then have
$\mathbbm{E}\left[n_{J_{k}}(\lnot\overline{a})\mid\mathrm{CE}\right]\leq\frac{1}{(1-\varepsilon^{\prime})^{K}\Delta_{k}}\left(R(T)+\overline{\theta}S_{k}\right)=:B_{k}=o(T).$
Therefore, we have
$\mathbbm{E}\left[n_{J_{k}}(\overline{a})\mid\mathrm{CE}\right]$ $=T-S_{k}-\mathbbm{E}\left[n_{J_{k}}(\lnot\overline{a})\mid\mathrm{CE}\right]$ (71)
$\geq T-(S_{k}+B_{k})$ (72)
$=T-o(T).$ (73)
Input: $T\in\mathbb{Z}_{>0},\varepsilon^{\prime}>0,R(T),\overline{a}\in\mathcal{A}\setminus\{a_{0}\},(\Delta_{k})_{k},\Theta=\{\theta_{1},\dots,\theta_{K}\}$
1 For each $k\in\{2,\dots,K\}$, define
$T_{k}\coloneqq\frac{2}{\varepsilon^{\prime}}\frac{1}{(1-\varepsilon^{\prime})^{K-k}}\left(R(T)+\sum_{s=k+1}^{K}\delta_{s}T_{s}\right),\quad\delta_{k}\coloneqq\frac{\theta_{k}-\theta_{k-1}}{2}.$
2Initialize $t\leftarrow 0$, $k\leftarrow K$;
3
4while $k\geq 2$ do
       // Exploration: Phase $k$ consists of $T_{k}$ periods
5       for $s=1,\dots,T_{k}$ do
6             $t\leftarrow t+1$;
7             $x_{t}(a)\leftarrow\mathbbm{1}\{a=\overline{a}\}$;
8             $p_{t}(a)\leftarrow(\theta_{k}-\delta_{k})\mathbbm{1}\{a=\overline{a}\}$;
9             observe $a_{t}$;
10
11      if $n_{I_{k}}(\overline{a})\coloneqq\sum_{t\in I_{k}}\mathbbm{1}\{a_{t}=\overline{a}\}\geq T_{k}/2$ then
             break; // Conclude $\theta=\theta_{k}$
12
13      else
14             $k\leftarrow k-1$;
15
16
17
18while $t\leq T$ do
       // Exploitation: price $\theta_{k}-\Delta_{k}\approx\theta_{k}$ is charged for $\overline{a}$ for the remaining periods
19       $t\leftarrow t+1$;
20       $x_{t}(a)\leftarrow\mathbbm{1}\{a=\overline{a}\}$;
21       $p_{t}(a)\leftarrow(\theta_{k}-\Delta_{k})\mathbbm{1}\{a=\overline{a}\}$;
22
Algorithm 5 Principal’s strategy against a no-ER learner
B.5 Proof of Lemma 2
Since $\overline{c}_{s}+\overline{\rho}_{s}\geq B$ with $s\coloneqq T_{0}$, we have
$\overline{c}_{T_{0}}\geq B-\sqrt{\frac{2\log T}{T_{0}}}.$
Since $\overline{c}_{T_{0}}\leq B-\Delta$, we have
$B-\sqrt{\frac{2\log T}{T_{0}}}\leq\overline{c}_{T_{0}}\leq B-\Delta,$
and thus
$\Delta\leq\sqrt{\frac{2\log T}{T_{0}}},$
which is equivalent to (6).
Input: Action set $\mathcal{A}$, time horizon $T$, agent’s type $\theta$, default action $a_{0}$, exploration length $T_{1}$
1 Phase 1: Exploration
2 for $t\in[|\mathcal{A}|T_{1}]$ do
3       Choose $a_{t}\coloneqq a_{k}$, where $k\coloneqq t\ (\mathrm{mod}\ |\mathcal{A}|)$
4       Observe allocation $x_{t}(a_{t})$, and payment $p_{t}(a_{t})$
5 end for
6Choose action $a^{*}\in\mathcal{A}$ with the highest empirical reward, i.e.,
$a^{*}\in\mathop{\rm arg~{}max}\limits_{a\in\mathcal{A}}\frac{1}{T_{1}}\sum_{t=1}^{|\mathcal{A}|T_{1}}\mathbbm{1}\{a_{t}=a\}\left(\theta x_{t}(a_{t})-p_{t}(a_{t})\right).$
Record the lower confidence bounds
$\mathrm{LCB}_{x}^{1}\coloneqq\left[\frac{1}{T_{1}}\sum_{t=1}^{|\mathcal{A}|T_{1}}\mathbbm{1}\{a_{t}=a^{*}\}x_{t}(a_{t})\right]-\overline{\rho}_{T_{1}},\quad\mathrm{UCB}_{p}^{1}\coloneqq\left[\frac{1}{T_{1}}\sum_{t=1}^{|\mathcal{A}|T_{1}}\mathbbm{1}\{a_{t}=a^{*}\}p_{t}(a_{t})\right]+\overline{\rho}_{T_{1}},$
where
$\overline{\rho}_{s}\coloneqq\sqrt{\frac{2\log T}{s}}\quad(s\in\mathbb{Z}_{>0}).$
7Phase 2: Exploitation with Protection
8 Initialize $s\coloneqq 1$
9 while True do
10       Play action $a^{*}$
11       Observe $x_{t}(a^{*})$ and $p_{t}(a^{*})$
12       Compute upper confidence intervals:
$\mathrm{UCB}_{x}^{2}(s)\coloneqq\sum_{t=|\mathcal{A}|T_{1}+1}^{|\mathcal{A}|T_{1}+s}x_{t}(a^{*})+\overline{\rho}_{s},\quad\mathrm{LCB}_{p}^{2}(s)\coloneqq\sum_{t=|\mathcal{A}|T_{1}+1}^{|\mathcal{A}|T_{1}+s}x_{t}(a^{*})-\overline{\rho}_{s}$
if $\mathrm{UCB}_{x}^{2}(s)<\mathrm{LCB}_{x}^{1}$ or $\mathrm{LCB}_{p}^{2}(s)>\mathrm{UCB}_{p}^{1}$ then
13            break
14       end if
15      $s\leftarrow s+1$
16 end while
17Phase 3: Punishment
18 while True do
19       Play the default action $a_{0}$ for all remaining periods
20 end while
21
Algorithm 6 Explore-Exploit-Punish (EEP)
B.6 Proof of Theorem 2
Assume $T\geq T_{1}$.
Proof of no-WER:
Suppose that the environment is stationary, i.e., for each action $a\in\mathcal{A}$, there is a fixed distribution $\mathbbm{P}(a)$ over $[0,1]\times[0,1/2]$ such that $(x_{t}(a),p_{t}(a))\operatorname{\overset{\mathrm{i.i.d.}}{\sim}}\mathbbm{P}(a)$. Let $\overline{x}(a)\coloneqq\mathbbm{E}_{\mathbbm{P}(a)}[x_{t}(a)]$ and $\overline{p}(a)\coloneqq\mathbbm{E}_{\mathbbm{P}(a)}[p_{t}(a)]$.
If we do not enter Phase 3, then the behavior of EEP is the same as that of Uniform Exploration, which has no-WER. Thus, it suffices to show that the probability of Phase 3 being triggered is sufficiently low when the environment is stationary.
By construction of the confidence interval (see Appendix A.2 for basic properties of confidence intervals), when entering Phase 2, with probability at least $1-o(T)$, the confidence interval $[\mathrm{LCB}_{x}^{1},\mathrm{UCB}_{x}^{1}]$ built in Phase 1 includes $\overline{x}(a^{*})$, where
$\mathrm{UCB}_{x}^{1}\coloneqq\left[\frac{1}{T_{1}}\sum_{t=1}^{|\mathcal{A}|T_{1}}\mathbbm{1}\{a_{t}=a^{*}\}x_{t}(a_{t})\right]+\overline{\rho}_{T_{1}}.$
Conditional on this “clean event”, by the same argument, we can show that the confidence interval built in Phase 2 contains $\overline{x}(a^{*})$ with probability at least $1-o(T)$, and we do not enter Phase 3 in this case. The similar argument applies to $\overline{p}(a)$. Therefore, the difference in payoffs between Uniform Exploration and EEP is at most $o(T)$, which implies that EEP is also of no-WER.
Proof of equilibrium payoff: Let
$\overline{x}_{T_{1}}(a)\coloneqq\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}x_{t}(a),\quad\overline{p}_{T_{1}}(a)\coloneqq\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}p_{t}(a)\quad\left(a\in\mathcal{A}\setminus\{a_{0}\}\right).$
First, consider the principal’s optimal payoff after Phase 1. For notational simplicity, let $\overline{\rho}\coloneqq\overline{\rho}_{T_{1}}$. If action $a_{0}$ is chosen at the end of Phase 1, she cannot do anything; she simply gets payoff 0.
Suppose that action $\overline{a}\neq a_{0}$ is chosen at the end of Phase 1, and the principal chose $(\overline{x}_{T_{1}}(\overline{a}),\overline{p}_{T_{1}}(\overline{a}))$. Principal’s payoff from period $|\mathcal{A}|T_{1}+1$ on is $\sum_{t\in I_{2}}p_{t}(\overline{a})$. Suppose that $\sum_{t\in I_{2}}p_{t}(\overline{a})=T_{2}(\overline{p}_{T_{1}}(\overline{a})+\overline{\rho})+T_{2}\Delta$ for some $\Delta\in\mathbb{R}$ at the optimum. Since the principal can make $T_{2}=T-|\mathcal{A}|T_{1}$ by choosing $p_{t}(\overline{a})=\overline{p}_{T_{1}}(\overline{a})+\overline{\rho}$, we have $\Delta\geq 0$. By Lemma 2,181818 Apply the second case of Lemma 2 with $B\coloneqq\mathrm{UCB}_{1}^{p}(\overline{a})=\overline{p}_{T_{1}}(\overline{a})+\overline{\rho}$. when $\Delta>0$, we have
$T_{2}\leq\min\left\{\frac{2\log T}{\Delta^{2}},T-|\mathcal{A}|T_{1}\right\}.$
Note that the RHS of this inequality becomes $\frac{2\log T}{\Delta^{2}}$ iff $\Delta\geq\sqrt{\frac{2\log T}{T-|\mathcal{A}|T_{1}}}$.
First, consider the case where
$T_{2}<\min\left\{\frac{2\log T}{\Delta^{2}},T-|\mathcal{A}|T_{1}\right\}.$
This is clearly suboptimal for the principal since she can increase her payoff by increasing $\Delta$ without affecting $T_{2}$. By the same reason, $\frac{2\log T}{\Delta^{2}}>T-|\mathcal{A}|T_{1}$ can never be optimal. Therefore, under the principal’s optimal strategy, we have $T_{2}=\frac{2\log T}{\Delta^{2}}$ with $\Delta\geq\sqrt{\frac{2\log T}{T-|\mathcal{A}|T_{1}}}$.
Given $\Delta$, principal’s payoff is
$T_{2}(\overline{p}_{T_{1}}(\overline{a})+\overline{\rho})+T_{2}\Delta=\frac{2\log T}{\Delta^{2}}\left(\overline{p}_{T_{1}}(\overline{a})+\overline{\rho}+\Delta\right)=:h(\Delta).$
We have
$h^{\prime}(\Delta)=\frac{2\log T}{\Delta^{3}}\left(-\Delta-2(\overline{p}_{T_{1}}(\overline{a})+\overline{\rho})\right)<0.$
Thus, $\Delta^{*}\coloneqq\sqrt{\frac{2\log T}{T-|\mathcal{A}|T_{1}}}=o(1)$ is optimal, and we have $T_{2}=T-T_{1}|\mathcal{A}|$. Therefore, the principal’s optimal cumulative payoff during Phase 2 is
$(T-|\mathcal{A}|T_{1})\left(\overline{p}_{T_{1}}(\overline{a})+\overline{\rho}+\Delta^{*}\right).$ (74)
Lastly, we consider the principal’s optimal strategy during Phase 1. We will show that, for any sufficiently large $T$, it is optimal to choose $\overline{p}_{T_{1}}(\overline{a})=p(\pi)$ for some $\overline{a}$, which is chosen at the end of Phase 1.
Observe that $x_{t}(\overline{a})\equiv 1$ is strictly optimal for the principal if action $\overline{a}$ is chosen at the end of Phase 1 as it minimizes the probability that the agent chooses $a_{0}$ at the end of Phase 1 for any $\theta$ given $\overline{p}_{T_{1}}(\overline{a})$. Since the learner chooses the action with the “lowest price” at the end of Phase 1, it is optimal for the principal that for all $t\in I_{1}$, for some $p\in\Theta$,
$\overline{p}_{T_{1}}(a)=\begin{cases}\overline{p}&(a\in\mathcal{A}\setminus\{\overline{a},a_{0}\})\\ p&(a=\overline{a})\\ 0&(a=a_{0})\end{cases},$
where $\overline{p}\coloneqq 1/2$ is the upper bound of the range of $p_{t}$. By (74), the principal’s expected payoff under belief $\pi$ is
$T_{1}\left[\left(|\mathcal{A}|-2\right)\overline{p}+p\right]+\left(T-|\mathcal{A}|T_{1}\right)\Pr_{\pi}\left(\theta\geq p\right)\left(p+\overline{\rho}+\Delta^{*}\right)=Tp\cdot\Pr_{\pi}\left(\theta\geq p\right)+o(T).$ (75)
Thus, it is optimal for the principal to choose $p\coloneqq p(\pi)$ under prior $\pi$ when $T$ is sufficiently large.
Proof of partial safety: For sufficiently large $T$, by the construction of $T_{1}$ and the confidence intervals, the principal cannot simultaneously achieve (i) making agents with all possible types choose $\overline{a}\neq a_{0}$ at the end of Phase 1, and (ii) always extracting full surplus. This can be shown by the same logic as in the proof of equilibrium payoff: after “posting price $p$” during Phase 1, the principal can change the price during Phase 2 at most $\Delta^{*}=o(1)$ on average.
B.7 Proof of Theorem 3
Recall that EER has no-WER and is partially safe.
No-WER
Assume that the environment is stationary. Suppose that $a^{*}\neq a_{0}$ at the end of Phase 1. In this case, the behavior of ESEP is the same as EEP except for the signaling phase, which lasts for $o(T)$ periods. Thus, the difference in average payoffs is $o(1)$.
Next, suppose that $a^{*}=a_{0}$ at the end of Phase 1. Under EEP, the learner keep choosing $a_{0}$ for the rest of the game. The behavior of ESEP can be different from EEP since it first enters the signaling phase, and then enters Phase 3. However, it will take suboptimal actions at most $o(T)$ periods by the property of confidence interval (Lemma 2). Thus, the difference in average payoffs is $o(1)$ as well. Therefore, ESEP is also of no-WER.
Partial safety
To see ESEP is partially safe, observe that the environment cannot do better against ESEP than against EEP since any potentially profitable deviation by the environment can be detected during Phase 3 (Lemma 2).
Welfare efficiency
By construction of ESEP, if $a^{*}=a_{0}$ is chosen at the end of Phase 1, it is approximately optimal for the environment to choose $x_{t}(\overline{a})\coloneqq 1$ and $p_{t}(\overline{a})\coloneqq 0$ in Phase 3 after the signaling phase: any possible profitable deviations are detected by the confidence interval method and the environment do better only by $o(1)$. The optimal payoff of the environment when $a^{*}\neq a_{0}$ is approximately the same as the one against EEP. In both cases, the learner obtains the good, i.e., $x_{t}(a_{t})=1$, for $T-o(T)$ periods if the adaptive environment best responds. Therefore, ESEP is welfare efficient. ∎
B.8 Consumer surplus under ESEP
Suppose the environment is adaptive and holds a belief $\pi$ over the agent’s type $\theta$, which is drawn from $\pi$ [^19] A natural interpretation is that the environment (e.g., an auction platform) observes data from past interactions and can estimate the distribution of agent types. The adaptive environment (the principal) interacts with a population of agents who all employ the same online learning algorithm $\mathcal{L}$. Anticipating the algorithm $\mathcal{L}$, the principal aims to maximize her expected payoff. We ask: what are the average payoffs to the principal and the agents under this setting?
Let $F$ denote the cumulative distribution function of $\pi$, and consider the associated demand curve $1-F(p)$. The principal can always post the monopoly price $p(\pi)$ in every period. Given that the agent commits to a fixed strategy and the principal is Bayesian, her ex ante expected payoff under the best response (as $T$ becomes large) is at least
$\underline{\mathrm{PS}}(\pi)+o(1),\quad\text{where}\quad\underline{\mathrm{PS}}(\pi)\coloneqq p(\pi)(1-F(p(\pi))).$
Because the total surplus in any interaction is bounded above by $\theta$, the agents’ average payoff under any online learning algorithm cannot exceed $\theta-\underline{\mathrm{PS}}(\pi)+o(1)$. The ESEP algorithm with $\varepsilon_{p}\approx 0$ achieves this upper bound. This is because: (i) it is optimal for the principal to post the monopoly price $p(\pi)$ during the exploration phase (Phase 1); and (ii) when the posted price is too high for a given type (i.e., $\theta<p(\pi)$), the signaling phase ensures that trade occurs in the exploitation phase at price $\varepsilon_{p}$.